### 13 Mar 2026 v2

# RC-positivity, comparison theorems and prescribed Hermitian-Yang-Mills tensors I

Abstract.
In this paper, we solve the prescribed Hermitian-Yang-Mills tensor problem. Let $E$ be a holomorphic vector bundle over a compact Kähler manifold $(M,\omega_{g})$. Suppose that there exists a smooth Hermitian metric $h_{0}$ on $E$ such that the Hermitian-Yang-Mills tensor $\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}$ is positive definite. Then for any Hermitian positive definite tensor $P\in\Gamma\left(M,E^{*}\otimes\overline{E}^{*}\right)$, there exists a unique smooth Hermitian metric $h$ on $E$ such that
$\Lambda_{\omega_{g}}\sqrt{-1}R^{h}=P.$
As applications, we obtain quantitative Chern number inequalities applicable to both holomorphic vector bundles and Fano manifolds. The proof is based on a new comparison theorem for Hermitian-Yang-Mills tensors.
1. Introduction
Let $(M,\omega_{g})$ be a compact Kähler manifold. It is well-known that the class of Ricci curvature $\mathrm{Ric}(\omega_{g})$ represents the first Chern class $c_{1}(M)$ up to a factor of $2\pi$, i.e.,
$\left[\mathrm{Ric}(\omega_{g})\right]=2\pi c_{1}(M)\in H^{1,1}(M,{\mathbb{R}}).$
The converse–given a cohomology class representative, can one find a Kähler metric whose Ricci form matches it–is the celebrated content of Yau’s solution to the Calabi conjecture [Cal57, Yau78]. Specifically, for any closed real $(1,1)$ form $\Omega$ representing $2\pi c_{1}(M)$, there exists a unique Kähler metric $\omega\in[\omega_{g}]$ such that
(1.1) $\mathrm{Ric}(\omega)=\Omega.$
By using $\partial$$\overline{\partial}$}-lemma, this equation is reduced to a Monge-Ampère type equation
$\left(\omega_{g}+\sqrt{-1}{\partial}{\overline{\partial}}\varphi\right)^{n}=e^{F}\omega_{g}^{n}.$
The Calabi-Yau theorem has indeed been extended beyond Kähler manifolds to more general Hermitian manifolds, particularly through the work of G. Székelyhidi, V. Tosatti and B. Weinkove, among others. We refer to [TW10a, TW10b, STW17] and the references therein.
The Uhlenbeck-Yau theorem [UY86] (see also Donaldson [Don85, Don87]) asserts that for a stable holomorphic vector bundle $E$ over a compact Kähler manifold $(M,\omega_{g})$, there exists a unique Hermitian-Einstein metric $h$ on $E$ up to scaling, satisfying
(1.2) $\Lambda_{\omega_{g}}\sqrt{-1}R^{h}=\lambda_{0}\cdot h,$
where $R^{h}\in\Gamma(M,\Lambda^{1,1}T^{*}M\otimes E^{*}\otimes\overline{E}^{*})$ is the Chern curvature tensor of $(E,h)$ and
$\lambda_{0}=\frac{n\int_{M}c_{1}(E)\wedge\omega_{g}^{n-1}}{\mathrm{rank}(E)\int_{M}\omega_{g}^{n}}.$
In particular, the Chern number inequality holds
(1.3) $\int_{M}\left((r-1)c^{2}_{1}(E)-2rc_{2}(E)\right)\wedge\omega_{g}^{n-2}\leq 0.$
Together with results from [Kob82, Lub83], this establishes a precise correspondence between the algebro-geometric notion of stability and the existence of Hermitian-Einstein metrics. This was generalized to the setting of Hermitian manifolds through the work [LY86] of Jun Li and S.-T. Yau.
The main result of this paper is the following analogue of the Calabi-Yau theorem, addressing the prescribed Hermitian-Yang-Mills tensor problem for holomorphic vector bundles on compact Kähler manifolds.
Theorem 1.1.
Let $E$ be a holomorphic vector bundle over a compact Kähler manifold $(M,\omega_{g})$. Suppose that there exists a smooth Hermitian metric $h_{0}$ on $E$ such that the Hermitian-Yang-Mills tensor $\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}$ is positive definite. Then for any Hermitian positive definite tensor $P\in\Gamma(M,E^{*}\otimes\overline{E}^{*})$, there exists a unique smooth Hermitian metric $h$ on $E$ such that
(1.4) $\Lambda_{\omega_{g}}\sqrt{-1}R^{h}=P.$
By Serre duality, the corresponding conjugate equation of (1.4) admits a solution when $\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}$ is (quasi-)negative definite; see Remark 5.2 for further details. The method for solving equation (1.4) differs significantly from those for equations (1.1) and (1.2), and we provide a brief overview of the key ideas involved. The uniqueness part of Theorem 1.1 is established by the following comparison theorem for Hermitian-Yang-Mills tensors of Hermitian vector bundles:
Theorem 1.2.
Let $E$ be a holomorphic vector bundle over a compact Kähler manifold $(M,\omega_{g})$. If $h$ and $h_{0}$ are smooth Hermitian metrics on $E$ satisfying $\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}>0$ and
(1.5) $\Lambda_{\omega_{g}}\sqrt{-1}R^{h}\leq\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}$
as Hermitian tensors in $\Gamma(M,E^{*}\otimes\overline{E}^{*})$, then $h\leq h_{0}$.
Remark 1.3.
Note that the curvature constraint $\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}>0$ in Theorem 1.1 is closely related to the notions of RC-positivity and uniform RC-positivity introduced in [Yang18, Yang20]. A comparison theorem analogous to Theorem 1.2 for RC-positive vector bundles was also established in [XYY24+]. This curvature constraint should be replaced by RC-positivity, and this replacement is carried out in a separate project on general Hermitian manifolds.
To establish the existence result, we first introduce two key function spaces: let $\mathrm{Herm}(E)$ denote the space of smooth Hermitian tensors on $E$, and $\mathrm{Herm}^{+}(E)$ be the space of smooth Hermitian metrics on $E$. We then define a Hermitian-Yang-Mills map
$G:\mathrm{Herm}^{+}(E)\rightarrow\mathrm{Herm}(E)$
by the formula
(1.6) $G(h)=\Lambda_{\omega_{g}}\sqrt{-1}R^{h}.$
We show that if there exists a smooth Hermitian metric $h_{0}$ such that $\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}>0$, then $G:G^{-1}\left(\mathrm{Herm}^{+}(E)\right)\rightarrow\mathrm{Herm}^{+}(E)$ is a bijection. The injectivity is given by Theorem 1.2. For the surjectivity, we show that the image of $G$
(1.7) $\mathscr{R}=\{P\in{\mathrm{Herm}^{+}(E)}\ |\ G(h)=\Lambda_{\omega_{g}}\sqrt{-1}R^{h}=P\text{ for some}\ h\in{\mathrm{Herm}^{+}(E)}\}$
is both open and closed in $\mathrm{Herm}^{+}(E)$ with respect to the $C^{\infty}$-topology. The openness of $\mathscr{R}$ (Theorem 5.6) is established via the implicit function theorem, leveraging the bijectivity of the linearization of $G$ (up to isomorphisms) at a reference metric $h_{1}$:
(1.8) $\mathscr{L}(\Psi)=\Delta_{\partial^{h_{1}}}\Psi+\Omega_{1}\cdot\Psi,$
where the positive definiteness of the Hermitian-Yang-Mills tensor $\Omega_{1}$ of $h_{1}$ ensures that $\mathscr{L}$ is an invertible self-adjoint elliptic operator. The closeness of $\mathscr{R}$ (Theorem 5.3) follows from two key analytic estimates: Theorem 1.2, which provides a uniform $C^{0}$-bound on metrics $h$ satisfying $G(h)=P$ for $P\in\mathscr{R}$, and a higher-order regularity estimate (Theorem 4.1), which controls the $C^{1}$-norm of such metrics. Together, these estimates imply that any sequence $\{P_{k}\}\subset\mathscr{R}$ converging in the $C^{\infty}$-topology has a preimage sequence $\{h_{k}\}$ that is precompact in the $C^{1}$-norm. By using $W^{k,p}$ and $C^{k,\alpha}$-estimates for elliptic equations, we obtain a smooth limit metric $h$ of some subsequence $\{h_{k_{s}}\}$ with $G(h)=\lim P_{k_{s}}=P$. Hence $\mathscr{R}$ is closed.
The Calabi-Yau theorem asserts that on a Fano manifold $M$, there exist Kähler metrics $\omega$ in each Kähler class satisfying $\mathrm{Ric}(\omega)>0$. In particular, $\Lambda_{\omega_{g}}\sqrt{-1}R^{g}>0$. As an immediate application of Theorem 1.1, we obtain the following result on the existence of Hermitian metrics with prescribed Hermitian-Yang-Mills tensors on Fano manifolds:
Corollary 1.4.
Let $M$ be a Fano manifold. Suppose that $P\in\Gamma(M,T^{*1,0}M\otimes T^{*0,1}M)$ is a Hermitian positive definite section. Then for any Kähler class $\omega]$, there exist a Kähler metric $\omega_{g}\in[\omega]$, and a unique Hermitian metric $h$ on $T^{1,0}M$ such that
(1.9) $\Lambda_{\omega_{g}}\sqrt{-1}R^{h}=P.$
The following special case is particularly worth highlighting:
Corollary 1.5.
On a Fano manifold $M$, for any Kähler metric $\omega_{g}$ with $\mathrm{Ric}(\omega_{g})>0$, there exists a unique Hermitian metric $h$ on $T^{1,0}M$ such that
(1.10) $\Lambda_{\omega_{g}}\sqrt{-1}R^{h}=g.$
Moreover, if $\mathrm{Ric}(\omega_{g})=c_{0}^{-1}\omega_{g}$ for some $c_{0}>0$, then $h=c_{0}g$.
By combining Theorem 1.1 and Theorem 1.2, we derive a quantitative version of Chern number inequalities, extending those associated with Hermitian-Einstein metrics.
Theorem 1.6.
Let $(M,\omega_{g})$ be a compact Kähler manifold and $E$ be a holomorphic vector bundle of rank $r$ over $M$. Suppose that there exists a Hermitian metric $h_{0}$ on $E$ satisfying
(1.11) $a\cdot h_{0}\leq\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}\leq b\cdot h_{0},$
for some constants $a,b\in{\mathbb{R}}$. Then the following Chern number inequality holds
(1.12) $\int_{M}\left((r-1)c^{2}_{1}(E)-2rc_{2}(E)\right)\wedge\omega_{g}^{n-2}\leq\frac{r(r-1)\left(b-a\right)^{2}}{8\pi^{2}n^{2}}\int_{M}\omega_{g}^{n},$
if $ab>0$ or $M$ is projective.
Remark 1.7.
It is evident that Theorem 1.1 bears profound theoretical connections to the Calabi-Yau theory, the Donaldson-Uhlenbeck-Yau theory, and the Kazdan-Warner theory. A number of related research initiatives are currently underway:
  1. $\bullet$
    A Higgs bundle-theoretic formulation of the main theorems will be presented in a separate work.
  2. $\bullet$
    A flow-based framework, constructed via variations of energy functionals, will be elaborated in subsequent sections.
  3. $\bullet$
    A refined notion of $\varepsilon$-stability, closely tied to the criterion in (1.11), is under development.
  4. $\bullet$
    Theorem 1.1 and Theorem 1.2 together provide an alternative proof of the Donaldson-Uhlenbeck-Yau theorem.
Acknowledgements. The second named author would like to thank Bing-Long Chen, Jixiang Fu, Kefeng Liu, Linlin Sun, Song Sun, Valentino Tosatti, Weiping Zhang and Xiping Zhu for inspired discussions.
2. Background materials
Let $(M,\omega_{g})$ be a compact Kähler manifold of dimension $n$ and $E$ be a holomorphic vector bundle over $M$ of rank $r$. Fix a smooth Hermitian metric $h_{0}$ on $E$, the Chern connection of $(E,h_{0})$ is denoted by $\nabla^{h_{0}}$. We shall use the natural decomposition
$\nabla^{h_{0}}={\partial}^{h_{0}}+{\overline{\partial}},$
where ${\partial}^{h_{0}}$ is the $(1,0)$ part and $\overline{\partial}$} is the $(0,1)$ part. The formal adjoints of ${\partial}^{h_{0}}$ and $\overline{\partial}$} with respect to $\omega_{g}$ and $h_{0}$ are denoted by ${\partial}^{*,h_{0}}$ and $\overline{{\partial}}{}^{*,h_{0}}$ respectively. Let $R^{h_{0}}$ be the Chern curvature tensor of $(E,h_{0})$. In local holomorphic coordinates $\{z^{i}\}$ of $M$ and local holomorphic basis $\{e_{\alpha}\}$ of $E$, the Chern curvature tensor $R^{h_{0}}$ can be written as
(2.1) $R^{h_{0}}=(R^{h_{0}})_{i\overline{j}\alpha}{}^{\beta}dz^{i}\wedge d\overline{z}^{j}\otimes e^{\alpha}\otimes e_{\beta}\in\Gamma(M,\Lambda^{1,1}T^{*}M\otimes E^{*}\otimes E).$
  1. (1)
    The (first) Chern-Ricci curvature of $(E,h_{0})$ is
    (2.2) $\mathrm{Ric}(h_{0})=\sqrt{-1}R^{h_{0}}_{i\overline{j}}dz^{i}\wedge d\overline{z}^{j}=\sqrt{-1}\sum_{\alpha}(R^{h_{0}})_{i\overline{j}\alpha}^{\alpha}dz^{i}\wedge d\overline{z}^{j}\in\Gamma(M,\Lambda^{1,1}T^{*}M).$
    It is well-known that the Chern-Ricci curvature represents the first Chern class $c_{1}(E)$ up to a factor of $2\pi$.
  2. (2)
    The Hermitian-Yang-Mills tensor $S^{h_{0}}=(S^{h_{0}})_{\alpha}^{\beta}e^{\alpha}\otimes e_{\beta}\in\Gamma(M,E^{*}\otimes E)$ of $(E,h_{0})$ is defined as
    (2.3) $S^{h_{0}}:=\Lambda_{\omega_{g}}\left(\sqrt{-1}R^{h_{0}}\right)=\left(g^{i\overline{j}}(R^{h_{0}})_{i\overline{j}\alpha}{}^{\beta}\right)e^{\alpha}\otimes e_{\beta}.$
    When no confusion arises, we also use the notation
    (2.4) $S^{h_{0}}=(S^{h_{0}})_{\alpha\overline{\beta}}e^{\alpha}\otimes\overline{e}^{\beta}\in\Gamma(M,E^{*}\otimes\overline{E}^{*})$
    where tensor components are
    (2.5) $S^{h_{0}}_{\alpha\overline{\beta}}=(S^{h_{0}})_{\alpha}^{\gamma}h_{0,\gamma\overline{\beta}}=g^{i\overline{j}}R^{h_{0}}_{i\overline{j}\alpha\overline{\beta}},\quad\mbox{and}\quad R^{h_{0}}_{i\overline{j}\alpha\overline{\beta}}=h_{0,\gamma\overline{\beta}}(R^{h_{0}})_{i\overline{j}\alpha}{}^{\gamma}$.
    We say $S^{h_{0}}>0$ if the Hermitian matrix $\left(S^{h_{0}}_{\alpha\overline{\beta}}\right)$ is positive definite over $M$.
  3. (3)
    The Chern scalar curvature of $(E,h_{0})$ is
    (2.6) $s_{h_{0}}=g^{i\overline{j}}R^{h_{0}}_{i\overline{j}}=\sum_{\alpha}(S^{h_{0}})_{\alpha}^{\alpha}=\sum h_{0}^{\alpha\overline{\beta}}S^{h_{0}}_{\alpha\overline{\beta}}=g^{i\overline{j}}h_{0}^{\alpha\overline{\beta}}R^{h_{0}}_{i\overline{j}\alpha\overline{\beta}}.$
Let $h\in\Gamma(M,E^{*}\otimes{\overline{E}}^{*})$ be another smooth Hermitian metric on $E$. Let $R^{h}$ and $S^{h}$ be the corresponding Chern curvature tensor and Hermitian-Yang-Mills tensor of $(E,h)$ respectively. There is an induced smooth section $H\in\Gamma(M,E^{*}\otimes E)$ given by
(2.7) $H=h\cdot h_{0}^{-1}.$
In local holomorphic basis $\{e_{\alpha}\}$ of $E$, $H=H_{\alpha}^{\beta}e^{\alpha}\otimes e_{\beta}$ is represented by
(2.8) $H_{\alpha}^{\beta}=h_{\alpha\overline{\gamma}}h_{0}^{\beta\overline{\gamma}}.$
It is clear that the inverse $H^{-1}=(H^{-1})_{\alpha}^{\beta}e^{\alpha}\otimes e_{\beta}\in\Gamma(M,E^{*}\otimes E)$ has components
(2.9) $(H^{-1})_{\alpha}^{\beta}=h_{0,\alpha\overline{\gamma}}h^{\beta\overline{\gamma}}.$
For $P=P_{\alpha}^{\beta}e^{\alpha}\otimes e_{\beta}\in\Omega^{p}(M,E^{*}\otimes E)$ and $Q=Q_{\alpha}^{\beta}e^{\alpha}\otimes e_{\beta}\in\Omega^{q}(M,E^{*}\otimes E)$, the product
(2.10) $P\cdot Q\in\Omega^{p+q}(M,E^{*}\otimes E)$
is defined as
(2.11) $P\cdot Q=(P_{\alpha}^{\beta}\wedge Q_{\beta}^{\gamma})\;e^{\alpha}\otimes e_{\gamma}.$
In particular, one has
(2.12) $H\cdot H^{-1}=H^{-1}\cdot H=\mathrm{Id}_{E}\in\Gamma(M,E^{*}\otimes E).$
The tensor $H$ will play a central role in what follows; we now present some of its key linear algebraic properties.
Definition 2.1.
A section $P\in\Gamma(M,E^{*}\otimes E)$ is called $h_{0}$-Hermitian if
(2.13) $h_{0}(Pv,w)=h_{0}(v,Pw),$
for any $v,w\in\Gamma(M,E)$. It is easy to see that $P=P_{\alpha}^{\beta}e^{\alpha}\otimes e_{\beta}$ is $h_{0}$-Hermitian if and only if
(2.14) $h_{0,\alpha\overline{\delta}}h_{0}^{\beta\overline{\gamma}}\overline{P}{}_{\gamma}^{\delta}=P_{\alpha}^{\beta}.$
An $h_{0}$-Hermitian section $P=P_{\alpha}^{\beta}e^{\alpha}\otimes e_{\beta}\in\Gamma(M,E^{*}\otimes E)$ is called positive definite if
$h_{0}(Pv,v)>0$
for all $v\neq 0$. This is equivalent to the matrix $\left(P_{\alpha\overline{\beta}}\right)=\left(h_{0,\gamma\overline{\beta}}P_{\alpha}^{\gamma}\right)$ being Hermitian positive definite. We also write $P>0$. Similarly, $P\geq 0$ is defined by requiring $h_{0}(Pv,v)\geq 0$.
Lemma 2.2.
  1. (1)
    $H=h\cdot h_{0}^{-1}\in\Gamma(M,E^{*}\otimes E)$ is $h_{0}$-Hermitian.
  2. (2)
    If $P\in\Gamma(M,E^{*}\otimes E)$ is $h_{0}$-Hermitian, for any $A,B\in\Gamma(M,E^{*}\otimes E)$,
    (2.15) $h_{0}(P\cdot A,B)=h_{0}(A,P\cdot B),\quad h_{0}(A\cdot P,B)=h_{0}(A,B\cdot P).$
  3. (3)
    If $A,B\in\Gamma(M,E^{*}\otimes E)$ are $h_{0}$-Hermitian and $A\geq 0,B\geq 0$, then
    (2.16) $h_{0}(A,B)\geq 0.$
    Moreover, for any section $C\in\Gamma(M,E^{*}\otimes E)$, one has
    (2.17) $h_{0}(A\cdot C\cdot B,C)\geq 0.$
Proof.
(1) follows from a straightforward computation that
(2.18) $h_{0,\alpha\overline{\delta}}h_{0}^{\beta\overline{\gamma}}\overline{H}{}_{\gamma}^{\delta}=h_{0,\alpha\overline{\delta}}h_{0}^{\beta\overline{\gamma}}h_{0}^{p\overline{\delta}}h_{p\overline{\gamma}}=H_{\alpha}^{\beta}.$
For (2), one has
$h_{0}(A\cdot P,B)$ $=$ $h_{0,\gamma\overline{q}}h_{0}^{\alpha\overline{p}}A_{\alpha}^{\beta}P_{\beta}^{\gamma}\overline{B}{}_{p}^{q}=h_{0,\gamma\overline{q}}h_{0}^{\alpha\overline{p}}A_{\alpha}^{\beta}\left(h_{0,\beta\overline{s}}h_{0}^{\gamma\overline{t}}\overline{P}{}_{t}^{s}\right)\overline{B}{}_{p}^{q}$
$=$ $h_{0,\beta\overline{s}}h_{0}^{\alpha\overline{p}}A_{\alpha}^{\beta}\overline{B}{}_{p}^{q}\overline{P}{}_{q}^{s}=h_{0}(A,B\cdot P),$
where the second identity holds since $P$ is $h_{0}$-Hermitian. The relation $h_{0}(P\cdot A,B)=h_{0}(A,P\cdot B)$ can be verified in a similar way.
For (3), since $A\geq 0$ and $B\geq 0$, $A^{1/2}$ and $B^{1/2}$ are well-defined and it is easy to deduce that both of them are $h_{0}$-Hermitian. Hence,
(2.19) $h_{0}(A,B)=h_{0}\left(A^{1/2},A^{1/2}B\right)=h_{0}\left(A^{1/2}B^{1/2},A^{1/2}B^{1/2}\right)\geq 0.$
A simple calculation shows that $h_{0}(A\cdot C\cdot B,C)=h_{0}(A^{1/2}\cdot C\cdot B^{1/2},A^{1/2}\cdot C\cdot B^{1/2})\geq 0$. ∎
3. A comparison theorem for Hermitian-Yang-Mills tensors
In this section, we present a proof of Theorem 1.2, which provides a comparison result for Hermitian-Yang-Mills tensors of abstract vector bundles. We then explore its applications, including the uniqueness of solutions to the prescribed Hermitian-Yang-Mills tensor equation and a version of $C^{2}$-estimate for Hermitian-Yang-Mills tensors.
Theorem 3.1.
Let $E$ be a holomorphic vector bundle over a compact Kähler manifold $(M,\omega_{g})$. If $h$ and $h_{0}$ are smooth Hermitian metrics on $E$ satisfying $\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}>0$ and
(3.1) $\Lambda_{\omega_{g}}\sqrt{-1}R^{h}\leq\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}$
as Hermitian tensors in $\Gamma(M,E^{*}\otimes\overline{E}^{*})$, then $h\leq h_{0}$.
Remark 3.2.
It is established via an analogous approach that: for an indecomposable vector bundle $E$, if the inequality
(3.2) $\Lambda_{\omega_{g}}\sqrt{-1}R^{h}\leq\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}$
as Hermitian tensors in $\Gamma(M,E^{*}\otimes E)$ with respect to $h_{0}$, then $h=ch_{0}$ for some $c\in\mathbb{R}^{+}$.
As an application, one has
Corollary 3.3.
Let $E$ be a holomorphic vector bundle over a compact Kähler manifold $(M,\omega_{g})$. Suppose that $h$ and $h_{0}$ are smooth Hermitian metrics on $E$. If
(3.3) $\Lambda_{\omega_{g}}\sqrt{-1}R^{h}=\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}>0$
as Hermitian tensors in $\Gamma(M,E^{*}\otimes\overline{E}^{*})$, then $h=h_{0}$.
In particular, we establish the uniqueness in Theorem 1.1:
Corollary 3.4.
Let $E$ be a holomorphic vector bundle over a compact Kähler manifold $(M,\omega_{g})$. Then for any Hermitian positive definite section $P\in\Gamma(M,E^{*}\otimes\overline{E}^{*})$, there exists at most one smooth Hermitian metric $h$ on $E$ such that
(3.4) $\Lambda_{\omega_{g}}\sqrt{-1}R^{h}=P.$
By rescaling, we obtain the following result, which is analogous to the $C^{2}$-estimate in the proof of the Calabi-Yau Theorem.
Corollary 3.5.
Suppose that $h$ and $h_{0}$ are two smooth Hermitian metrics on $E$. If $\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}>0$ and
(3.5) $\Lambda_{\omega_{g}}\sqrt{-1}R^{h}\leq\lambda\cdot\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}$
as Hermitian tensors in $\Gamma(M,E^{*}\otimes\overline{E}^{*})$ and $\lambda\in{\mathbb{R}}^{+}$, then $h\leq\lambda\cdot h_{0}$.
The proof of Theorem 3.1 starts with the following relation concerning Hermitian-Yang-Mills tensors.
Proposition 3.6.
Let $E$ be a holomorphic vector bundle over a compact Kähler manifold $(M,\omega_{g})$. Suppose that $h$ and $h_{0}$ are two smooth Hermitian metrics on $E$. Then
(3.6) $\Lambda_{\omega_{g}}\sqrt{-1}R^{h}-\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}={\partial}^{*,h_{0}}\left(({\partial}^{h_{0}}H)\cdot H^{-1}\right),$
as tensors in $\Gamma(M,E^{*}\otimes E)$, where $H=h\cdot h_{0}^{-1}$ and ${\partial}^{h_{0}}$ is the $(1,0)$ component of the Chern connection on $E^{*}\otimes E$ induced by $(E,h_{0})$.
Proof.
The Chern connection $\nabla^{h_{0}}$ is given by
(3.7) $\nabla_{i}^{h_{0}}e_{\beta}=(\Gamma_{0})_{i\beta}^{\mu}e_{\mu},\quad\nabla_{i}^{h_{0}}e^{\alpha}=-(\Gamma_{0})_{i\beta}^{\alpha}e^{\beta}.$
where $(\Gamma_{0})_{i\alpha}^{\beta}$ is the Christoffel symbol of the Chern connection on $(E,h_{0})$. Hence,
$\nabla_{i}^{h_{0}}H$ $=$ $\nabla^{h_{0}}_{i}\left(h_{0}^{\beta\overline{\gamma}}h_{\alpha\overline{\gamma}}\;e^{\alpha}\otimes e_{\beta}\right)$
$=$ ${\partial}_{i}\left(h_{0}^{\beta\overline{\gamma}}h_{\alpha\overline{\gamma}}\right)\;e^{\alpha}\otimes e_{\beta}+h_{0}^{\beta\overline{\gamma}}h_{\alpha\overline{\gamma}}\left(\nabla_{i}^{h_{0}}e^{\alpha}\right)\otimes e_{\beta}+h_{0}^{\beta\overline{\gamma}}h_{\alpha\overline{\gamma}}e^{\alpha}\otimes\left(\nabla_{i}^{h_{0}}e_{\beta}\right)$
$=$ $\left(\left({\partial}_{i}h_{0}^{\beta\overline{\gamma}}\right)h_{\alpha\overline{\gamma}}+h_{0}^{\beta\overline{\gamma}}({\partial}_{i}h_{\alpha\overline{\gamma}})-h_{0}^{\beta\overline{\gamma}}h_{\mu\overline{\gamma}}(\Gamma_{0})_{i\alpha}^{\mu}+h_{0}^{\mu\overline{\gamma}}h_{\alpha\overline{\gamma}}(\Gamma_{0})_{i\mu}^{\beta}\right)\;e^{\alpha}\otimes e_{\beta}$
$=$ $h_{0}^{\beta\overline{\gamma}}\left(({\partial}_{i}h_{\alpha\overline{\gamma}})-h_{\mu\overline{\gamma}}(\Gamma_{0})_{i\alpha}^{\mu}\right)\;e^{\alpha}\otimes e_{\beta}.$
Therefore,
(3.8) ${\partial}^{h_{0}}H=h_{0}^{\beta\overline{\gamma}}\left(({\partial}_{i}h_{\alpha\overline{\gamma}})-h_{\mu\overline{\gamma}}(\Gamma_{0})_{i\alpha}^{\mu}\right)\;(dz^{i}\otimes e^{\alpha}\otimes e_{\beta})$
and
(3.9) $({\partial}^{h_{0}}H)\cdot H^{-1}$ $=$ $h^{\beta\overline{\gamma}}\left(({\partial}_{i}h_{\alpha\overline{\gamma}})-h_{\mu\overline{\gamma}}(\Gamma_{0})_{i\alpha}^{\mu}\right)\;(dz^{i}\otimes e^{\alpha}\otimes e_{\beta})$
$=$ $\left(\Gamma_{i\alpha}^{\beta}-(\Gamma_{0})_{i\alpha}^{\beta}\right)\;(dz^{i}\otimes e^{\alpha}\otimes e_{\beta})$
where $\Gamma_{i\alpha}^{\beta}$ is the Christoffel symbol of the Chern connection $\nabla^{h}$ of $h$. Moreover, the Chern curvature components are given by
(3.10) $(R^{h})_{i\overline{j}\alpha}{}^{\beta}=-\frac{{\partial}\Gamma_{i\alpha}^{\beta}}{{\partial}\overline{z}^{j}},\quad(R^{h_{0}})_{i\overline{j}\alpha}{}^{\beta}=-\frac{{\partial}(\Gamma_{0})_{i\alpha}^{\beta}}{{\partial}\overline{z}^{j}},$
and this implies the curvature identity
(3.11) $\overline{{\partial}}{}\left(({\partial}^{h_{0}}H)\cdot H^{-1}\right)=\left((R^{h})_{i\overline{j}\alpha}{}^{\beta}-(R^{h_{0}})_{i\overline{j}\alpha}{}^{\beta}\right)(dz^{i}\wedge d\overline{z}{}^{j}\otimes e^{\alpha}\otimes e_{\beta})=R^{h}-R^{h_{0}}.$
On the other hand, by using the Bochner-Kodaira formula
(3.12) ${\partial}^{*,h_{0}}=\sqrt{-1}[\Lambda_{\omega_{g}},\overline{{\partial}}],$
one obtains the difference (3.6) of the Hermitian-Yang-Mills tensors. ∎
Proof of Theorem 3.1.
To simplify notations, we set $\Omega=S^{h_{0}}\in\Gamma(M,E^{*}\otimes E)$ and define
(3.13) $\Phi:=h_{0}^{\alpha\overline{\gamma}}S^{h}_{\beta\overline{\gamma}}\;e^{\beta}\otimes e_{\alpha}\in\Gamma(M,E^{*}\otimes E).$
It’s easy to see that
(3.14) $\Phi=h_{0}^{\alpha\overline{\gamma}}h_{\delta\overline{\gamma}}(S^{h})^{\delta}_{\beta}\;e^{\beta}\otimes e_{\alpha}=S^{h}\cdot H.$
On the other hand, by (3.6),
(3.15) $S^{h}=S^{h_{0}}+{\partial}^{*,h_{0}}({\partial}^{h_{0}}H\cdot H^{-1}),$
and in particular, we obtain the identity
(3.16) $\Phi\cdot H^{-1}=\Omega+{\partial}^{*,h_{0}}({\partial}^{h_{0}}H\cdot H^{-1}).$
This equation can be rewritten in the following form:
(3.17) $\Omega\cdot(H^{-1}-\mathrm{Id}_{E})={\partial}^{*,h_{0}}({\partial}^{h_{0}}H\cdot H^{-1})+(\Omega-\Phi)\cdot H^{-1}.$
We claim that
(3.18) $H\leq\operatorname{Id}_{E}$
with respect to $h_{0}$ which also implies $h\leq h_{0}$. Indeed, for any $v=v^{\alpha}\cdot e_{\alpha}\in\Gamma(M,E)$,
(3.19) $h_{0}(Hv,v)=h_{0,\beta\overline{\gamma}}H^{\beta}_{\alpha}v^{\alpha}\overline{v}{}^{\gamma}=h_{0,\beta\overline{\gamma}}h_{0}^{\beta\overline{\delta}}h_{\alpha\overline{\delta}}v^{\alpha}\overline{v}{}^{\gamma}=h_{\alpha\overline{\gamma}}v^{\alpha}\overline{v}{}^{\gamma}=h(v,v).$
Hence, the relation $H\leq\operatorname{Id}_{E}$ implies that for all $v\in\Gamma(M,E)$
(3.20) $h(v,v)=h_{0}(Hv,v)\leq h_{0}(v,v).$
To prove (3.18), we define $\kappa:M\rightarrow\mathbb{R}$ as
(3.21) $\kappa(x):=\sup_{v\in E_{x},v\neq 0}\frac{h_{0}(Hv,v)}{h_{0}(v,v)}.$
The maximal eigenvalue of $H$ with respect to $h_{0}$ is
(3.22) $\Lambda:=\sup_{x\in M}\kappa(x).$
We argue by contradiction and assume that (3.18) does not hold. In this case, one has
(3.23) $\Lambda>1.$
For any real number $\varepsilon>0$, we consider the following test section
(3.24) $W_{\varepsilon}:=\left((\Lambda+\varepsilon)\mathrm{Id}_{E}-H\right)^{-2n}\cdot H\in\Gamma(M,E^{*}\otimes E).$
Since the maximal eigenvalue of $H$ is $\Lambda$,
(3.25) $(\Lambda+\varepsilon)\mathrm{Id}_{E}-H>0,$
and so
(3.26) $\left((\Lambda+\varepsilon)\mathrm{Id}_{E}-H\right)^{-2n}>0$
with respect to $h_{0}$. By part $(2)$ of Lemma 2.2, one has
$\left((\Omega-\Phi)\cdot H^{-1},W_{\varepsilon}\right)_{h_{0}}$ $=$ $\left((\Omega-\Phi)\cdot H^{-1},((\Lambda+\varepsilon)\mathrm{Id}_{E}-H)^{-2n}\cdot H\right)_{h_{0}}$
$=$ $\left(\Omega-\Phi,((\Lambda+\varepsilon)\mathrm{Id}_{E}-H)^{-2n}\right)_{h_{0}}.$
On the other hand, $\Phi$ is Hermitian with respect to $h_{0}$:
(3.27) $h_{0,\alpha\overline{\delta}}h_{0}^{\beta\overline{\gamma}}\overline{\Phi}{}_{\gamma}^{\delta}=h_{0,\alpha\overline{\delta}}h_{0}^{\beta\overline{\gamma}}h_{0}^{p\overline{\delta}}S^{h}_{p\overline{\gamma}}=h_{0}^{\beta\overline{\gamma}}S^{h}_{\alpha\overline{\gamma}}=\Phi_{\alpha}^{\beta}.$
Hence, by the curvature assumption $\left(S^{h}_{\alpha\overline{\beta}}\right)\leq\left(S^{h_{0}}_{\alpha\overline{\beta}}\right)$, we have
(3.28) $\Omega-\Phi\geq 0$
with respect to $h_{0}$. By part $(3)$ of Lemma 2.2,
(3.29) $\left((\Omega-\Phi)\cdot H^{-1},W_{\varepsilon}\right)_{h_{0}}=\left(\Omega-\Phi,((\Lambda+\varepsilon)\mathrm{Id}_{E}-H)^{-2n}\right)_{h_{0}}\geq 0.$
For any $k\geq 0$ and any fixed point $x\in M$, we assume that $g_{i\overline{j}}(x)=\delta_{ij}$ and so at point $x$
$\langle{\partial}^{h_{0}}H\cdot H^{-1},{\partial}^{h_{0}}H^{k}\rangle_{h_{0}}$ $=$ $\sum_{p=1}^{k}\langle{\partial}^{h_{0}}H\cdot H^{-1},H^{p-1}\cdot{\partial}^{h_{0}}H\cdot H^{k-p}\rangle_{h_{0}}$
$=$ $\sum_{i=1}^{n}\sum_{p=1}^{k}\langle H^{p-1}\cdot{\partial}_{i}^{h_{0}}H\cdot H^{k-p-1},{\partial}_{i}^{h_{0}}H\rangle_{h_{0}}$
$\geq$ $0,$
where the last inequality holds by part (3) of Lemma 2.2. Since $H$ has maximal eigenvalue $\Lambda$, one has the Taylor expansion
(3.30) $\left((\Lambda+\varepsilon)\mathrm{Id}_{E}-H\right)^{-2n}=\sum_{k=0}^{\infty}{2n+k-1\choose k}(\Lambda+\varepsilon)^{-(k+2n)}H^{k}.$
By using this expansion and integration by parts, one gets
$\left({\partial}^{*,h_{0}}({\partial}^{h_{0}}H\cdot H^{-1}),W_{\varepsilon}\right)_{h_{0}}$
$=$ $\left({\partial}^{h_{0}}H\cdot H^{-1},{\partial}^{h_{0}}\left(((\Lambda+\varepsilon)\mathrm{Id}_{E}-H)^{-2n}\cdot H\right)\right)_{h_{0}}$
$=$ $\sum_{k=0}^{\infty}{2n+k-1\choose k}(\Lambda+\varepsilon)^{-(k+2n)}\left({\partial}^{h_{0}}H\cdot H^{-1},{\partial}^{h_{0}}H^{k+1}\right)_{h_{0}}\geq 0.$
By pairing with $W_{\varepsilon}$ in (3.17), one gets
(3.31) $\left(\Omega\cdot(H^{-1}-\mathrm{Id}_{E}),W_{\varepsilon}\right)_{h_{0}}=\left({\partial}^{*,h_{0}}({\partial}^{h_{0}}H\cdot H^{-1}),W_{\varepsilon}\right)_{h_{0}}+\left((\Omega-\Phi)\cdot H^{-1},W_{\varepsilon}\right)_{h_{0}}\geq 0.$
In the following, we estimate $\left(\Omega\cdot(H^{-1}-\mathrm{Id}_{E}),W_{\varepsilon}\right)_{h_{0}}$ using a different method and derive a contradiction. Let $K$ and $k$ be the largest and smallest eigenvalues of $\Omega$ with respect to $h_{0}$ respectively, i.e.,
(3.32) $K=\sup_{x\in M}\sup_{v\in E_{x},v\neq 0}\frac{h_{0}(\Omega v,v)}{h_{0}(v,v)}>0,\quad k=\inf_{x\in M}\inf_{v\in E_{x},v\neq 0}\frac{h_{0}(\Omega v,v)}{h_{0}(v,v)}>0.$
In particular,
(3.33) $\mathrm{tr}_{E}\Omega\leq rK.$
where $r$ is the rank of $E$. We claim that
  1. (I)
    For any point $x\in M$, one obtains the uniform estimate
    (3.34) $\left\langle\Omega,((\Lambda+\varepsilon)\mathrm{Id}_{E}-H)^{-2n}\cdot(\mathrm{Id}_{E}-H)\right\rangle_{h_{0}}(x)\leq rK(\Lambda+\varepsilon-1)^{-2n}.$
  2. (II)
    For any $x\in M$ satisfying $\kappa(x)\geq(\Lambda+1)/2$, the following refined estimate holds:
    (3.35) $\left\langle\Omega,((\Lambda+\varepsilon)\mathrm{Id}_{E}-H)^{-2n}\cdot(\mathrm{Id}_{E}-H)\right\rangle_{h_{0}}(x)\leq rK(\Lambda+\varepsilon-1)^{-2n}-\frac{k(\Lambda-1)}{2(\Lambda+\varepsilon-\kappa(x))^{2n}}.$
To prove (I), let $f$ be an auxiliary function defined by
(3.36) $f(t)=\frac{1-t}{(\Lambda+\varepsilon-t)^{2n}},\quad t\in(0,\Lambda].$
One can see clearly that $f(H)$ is a well-defined smooth section in $\Gamma(M,E^{*}\otimes E)$ and it is $h_{0}$-Hermitian. Moreover,
(3.37) $\left\langle\Omega,((\Lambda+\varepsilon)\mathrm{Id}_{E}-H)^{-2n}\cdot(\mathrm{Id}_{E}-H)\right\rangle_{h_{0}}=\left\langle\Omega,f(H)\right\rangle_{h_{0}}=\mathrm{tr}_{E}(\Omega\cdot f(H)),$
where the last identity holds since both $\Omega$ and $f(H)$ are $h_{0}$-Hermitian. Let
$0<\lambda_{1}\leq\cdots\leq\lambda_{r}\leq\Lambda$
be eigenvalues of $H$ at point $x\in M$. The corresponding eigenvalues of $f(H)$ are
$f(\lambda_{1}),\cdots,f(\lambda_{r}).$
It is clear that for any $t\in(0,\Lambda]$
(3.38) $f(t)\leq(\Lambda+\varepsilon-1)^{-2n}.$
In particular, one has
(3.39) $f(H)\leq(\Lambda+\varepsilon-1)^{-2n}\mathrm{Id}_{E}.$
Therefore,
(3.40) $\mathrm{tr}_{E}(\Omega\cdot f(H))\leq(\Lambda+\varepsilon-1)^{-2n}\mathrm{tr}_{E}(\Omega)\leq rK(\Lambda+\varepsilon-1)^{-2n}.$
This establishes (3.34).
For (II), if $x\in M$ satisfies $\kappa(x)\geq(\Lambda+1)/2$, then
(3.41) $f(\lambda_{r})=f(\kappa(x))=\frac{1-\kappa(x)}{(\Lambda+\varepsilon-\kappa(x))^{2n}}\leq-\frac{\Lambda-1}{2(\Lambda+\varepsilon-\kappa(x))^{2n}}.$
Suppose that $v_{x}\in E_{x}$ is an eigenvector of $H$ with unit length corresponding to $\lambda_{r}=\kappa(x)$. By (3.39), (3.41) and the fact that $\Lambda>1$, one achieves the following estimate at point $x$
(3.42) $f(H)\leq(\Lambda+\varepsilon-1)^{-2n}\mathrm{Id}_{E}-\frac{\Lambda-1}{2(\Lambda+\varepsilon-\kappa(x))^{2n}}\;v_{x}^{*}\otimes v_{x}.$
In particular, one has
$\mathrm{tr}_{E}(\Omega\cdot f(H))$ $\leq$ $(\Lambda+\varepsilon-1)^{-2n}\mathrm{tr}_{E}(\Omega)-\frac{\Lambda-1}{2(\Lambda+\varepsilon-\kappa(x))^{2n}}\mathrm{tr}_{E}(\Omega\cdot(v_{x}^{*}\otimes v_{x}))$
$\leq$ $rK(\Lambda+\varepsilon-1)^{-2n}-\frac{k(\Lambda-1)}{2(\Lambda+\varepsilon-\kappa(x))^{2n}}.$
This establishes the inequality (3.35).
Now we fix a point $x_{0}\in M$ with $\kappa(x_{0})=\Lambda$. Since $\kappa(x)$ is a Lipschitz function with Lipschitz constant $L=L(H,M,\omega_{g},h_{0})>0$, we set
$\delta=\frac{\Lambda-1}{2L}>0.$
For any $x\in M$ with $d(x,x_{0})<\delta$, one has $|\kappa(x)-\kappa(x_{0})|\leq Ld(x,x_{0})$, and so
(3.43) $\kappa(x)\geq\kappa(x_{0})-Ld(x,x_{0})=\Lambda-Ld(x,x_{0})\geq\frac{\Lambda+1}{2}.$
By (3.35),
$\int_{B(x_{0},\delta)}\langle\Omega,((\Lambda+\varepsilon)\mathrm{Id}_{E}-H)^{-2n}\cdot(\mathrm{Id}_{E}-H)\rangle_{h_{0}}\cdot\frac{\omega_{g}^{n}}{n!}$
$\leq$ $\frac{rK}{(\Lambda+\varepsilon-1)^{2n}}\int_{B(x_{0},\delta)}\frac{\omega_{g}^{n}}{n!}-\int_{B(x_{0},\delta)}\frac{k(\Lambda-1)}{2(\Lambda+\varepsilon-\kappa(x))^{2n}}\cdot\frac{\omega_{g}^{n}}{n!}$
$\leq$ $\frac{rK}{(\Lambda+\varepsilon-1)^{2n}}\int_{B(x_{0},\delta)}\frac{\omega_{g}^{n}}{n!}-\int_{B(x_{0},\delta)}\frac{k(\Lambda-1)}{2(\varepsilon+Ld(x,x_{0}))^{2n}}\cdot\frac{\omega_{g}^{n}}{n!}$
since $\Lambda-\kappa(x)\leq Ld(x,x_{0})$. Moreover, by (3.34), one obtains
$\left(\Omega,((\Lambda+\varepsilon)\mathrm{Id}_{E}-H)^{-2n}\cdot(\mathrm{Id}_{E}-H)\right)_{h_{0}}$
$\leq$ $\frac{rK\mathrm{Vol}(M,\omega_{g})}{(\Lambda+\varepsilon-1)^{2n}}-\int_{B(x_{0},\delta)}\frac{k(\Lambda-1)}{2(\varepsilon+Ld(x,x_{0}))^{2n}}\cdot\frac{\omega_{g}^{n}}{n!}.$
On the other hand, we established in (3.31) that
(3.44) $\left(\Omega,((\Lambda+\varepsilon)\mathrm{Id}_{E}-H)^{-2n}\cdot(\mathrm{Id}_{E}-H)\right)_{h_{0}}=\left(\Omega\cdot(H^{-1}-\mathrm{Id}_{E}),W_{\varepsilon}\right)_{h_{0}}\geq 0.$
By the monotone convergence theorem, one has
$0$ $\leq$ $\liminf_{\varepsilon\rightarrow 0^{+}}\left(\Omega,((\Lambda+\varepsilon)\mathrm{Id}_{E}-H)^{-2n}\cdot(\mathrm{Id}_{E}-H)\right)_{h_{0}}$
$\leq$ $\frac{rK\mathrm{Vol}(M,\omega_{g})}{(\Lambda-1)^{2n}}-\frac{k(\Lambda-1)}{2L^{2n}}\int_{B(x_{0},\delta)}\frac{1}{d(x,x_{0})^{2n}}\cdot\frac{\omega_{g}^{n}}{n!}=-\infty.$
That’s a contradiction. Therefore, one deduces $\Lambda\leq 1$ and so $h\leq h_{0}$. ∎
4. Priori estimates
In this section, we establish priori $C^{1}$-estimates for $H=h\cdot h_{0}^{-1}$ where $h$ and $h_{0}$ are two smooth Hermitian metrics on $E$. Recall that
  1. (1)
    $S^{h}$ and $S^{h_{0}}$ are the Hermitian-Yang-Mills tensors of $h$ and $h_{0}$ respectively;
  2. (2)
    $S^{h_{0}}>0$ with respect to $h_{0}$;
  3. (3)
    We use notations
    (4.1) $H=h_{\alpha\overline{\gamma}}h_{0}^{\beta\overline{\gamma}}\;e^{\alpha}\otimes e_{\beta},\quad\Omega=(S^{h_{0}})_{\alpha}^{\beta}\;e^{\alpha}\otimes e_{\beta},\quad\Phi=h_{0}^{\beta\overline{\gamma}}S^{h}_{\alpha\overline{\gamma}}\;e^{\alpha}\otimes e_{\beta}.$
    Note that $S^{h_{0}}_{\alpha\overline{\gamma}}=(S^{h_{0}})_{\alpha}^{\delta}h_{0,\delta\overline{\gamma}}$ and $S^{h}_{\alpha\overline{\gamma}}=(S^{h})_{\alpha}^{\delta}h_{\delta\overline{\gamma}}$.
The main result of this section is following uniform $C^{1}$-estimate for $H$. By using this uniform estimate and an elliptic equation satisfied by $H$, we derive uniform higher-order $W^{k,p}$ and $C^{k,\alpha}$-estimates.
Theorem 4.1.
If there exists a constant $C_{1}\geq 1$ such that
(4.2) $C_{1}^{-1}\left(h_{0,\alpha\beta}\right)\leq\left(S^{h}_{\alpha\overline{\beta}}\right)\leq C_{1}\left(h_{0,\alpha\overline{\beta}}\right)$
as Hermitian matrices, then one has
(4.3) $|H|_{C^{1}(M,\omega_{g},h_{0})}\leq C_{16},$
where $C_{16}$ depends on $M,\omega_{g},C_{1},h_{0},r$ and $\|\Phi\|_{C^{1}(M,\omega_{g},h_{0})}$.
We begin with the following estimate.
Proposition 4.2.
If there exists a constant $C_{1}\geq 1$ such that
(4.4) $C_{1}^{-1}\left(h_{0,\alpha\beta}\right)\leq\left(S^{h}_{\alpha\overline{\beta}}\right)\leq C_{1}\left(h_{0,\alpha\overline{\beta}}\right)$
as Hermitian matrices, then there exists a constant $C_{2}=C_{2}(M,\omega_{g},h_{0},C_{1})\geq 1$ such that
(4.5) $C_{2}^{-1}h_{0}\leq h\leq C_{2}h_{0}.$
Proof.
Since $M$ is compact and $S^{h_{0}}>0$, there exists $C_{3}=C_{3}(M,\omega_{g},h_{0})>0$ such that
(4.6) $C_{3}^{-1}\left(h_{0,\alpha\beta}\right)\leq\left(S^{h_{0}}_{\alpha\overline{\beta}}\right)\leq C_{3}\left(h_{0,\alpha\overline{\beta}}\right).$
By the condition (4.4), one has
(4.7) $(C_{1}C_{3})^{-1}\left(S^{h_{0}}_{\alpha\overline{\beta}}\right)\leq\left(S^{h}_{\alpha\overline{\beta}}\right)\leq C_{1}C_{3}\left(S^{h_{0}}_{\alpha\overline{\beta}}\right),$
as Hermitian matrices. By Corollary 3.5, one has
$C_{2}^{-1}h_{0}\leq h\leq C_{2}h_{0}$
and $C_{2}=C_{1}C_{3}$. ∎
Proposition 4.3.
Suppose that there exists a constant $C_{4}\geq 1$ such that
(4.8) $C_{4}^{-1}h_{0}\leq h\leq C_{4}h_{0}.$
Then for any $A\in\Gamma(M,E^{*}\otimes E)$,
(4.9) $C^{-1}_{4}h_{0}(A\cdot H^{-1},A)\leq h(A\cdot H^{-1},A\cdot H^{-1})\leq C_{4}h_{0}(A\cdot H^{-1},A)$
and
(4.10) $C_{4}^{-2}h_{0}(A,A)\leq h(A\cdot H^{-1},A\cdot H^{-1})\leq C_{4}^{2}h_{0}(A,A).$
In particular,
(4.11) $C_{4}^{-2}h_{0}({\partial}^{h_{0}}H,{\partial}^{h_{0}}H)\leq h({\partial}^{h_{0}}H\cdot H^{-1},{\partial}^{h_{0}}H\cdot H^{-1})\leq C_{4}^{2}h_{0}({\partial}^{h_{0}}H,{\partial}^{h_{0}}H).$
Moreover,
(4.12) $\mathrm{tr}_{E}\left(\Lambda_{\omega_{g}}\left(\sqrt{-1}{\partial}^{h_{0}}H\cdot H^{-1}\cdot{\overline{\partial}}H\right)\right)\geq C_{4}^{-1}h\left({\partial}^{h_{0}}H\cdot H^{-1},{\partial}^{h_{0}}H\cdot H^{-1}\right).$
Proof.
For (4.9), we fix a point $x\in M$ and assume that $h_{0,\alpha\overline{\beta}}(x)=\delta_{\alpha\beta}$ and $h_{\alpha\overline{\beta}}(x)=\lambda_{\alpha}\delta_{\alpha\beta}$. In this case $H(x)=\sum\lambda_{\alpha}(x)e^{\alpha}\otimes e_{\alpha}$ and the condition $C_{4}^{-1}h_{0}\leq h\leq C_{4}h_{0}$ gives
(4.13) $C_{4}^{-1}\leq\lambda_{\mu}\leq C_{4}.$
If $A=A_{\alpha}^{\beta}e^{\alpha}\otimes e_{\beta}$, then
$A\cdot H^{-1}(x)=\sum_{\alpha,\beta}A_{\alpha}^{\beta}(x)\lambda^{-1}_{\beta}e^{\alpha}\otimes e_{\beta}$
and $h_{0}(A,A)=\sum_{\alpha,\beta}|A_{\alpha}^{\beta}|^{2}$,
(4.14) $h_{0}(A\cdot H^{-1},A)=\sum_{\alpha,\beta}|A_{\alpha}^{\beta}|^{2}\lambda^{-1}_{\beta},\quad h(A\cdot H^{-1},A\cdot H^{-1})=\sum_{\alpha,\beta}|A_{\alpha}^{\beta}|^{2}\lambda^{-1}_{\alpha}\lambda^{-1}_{\beta}.$
Therefore,
(4.15) $C^{-1}_{4}h_{0}(A\cdot H^{-1},A)\leq h(A\cdot H^{-1},A\cdot H^{-1})\leq C_{4}h_{0}(A\cdot H^{-1},A)$
and
(4.16) $C_{4}^{-2}h_{0}(A,A)\leq h(A\cdot H^{-1},A\cdot H^{-1})\leq C_{4}^{2}h_{0}(A,A).$
Moreover, at the point $x\in M$, we can also assume $g_{i\overline{j}}(x)=\delta_{ij}$. Hence
(4.17) $\mathrm{tr}_{E}\left(\Lambda_{\omega_{g}}\left(\sqrt{-1}{\partial}^{h_{0}}H\cdot H^{-1}\cdot{\overline{\partial}}H\right)\right)=\sum_{i=1}^{n}\mathrm{tr}_{E}\left(\nabla_{i}^{h_{0}}H\cdot H^{-1}\cdot\overline{{\partial}}_{i}H\right).$
On the other hand, one has
(4.18) $\mathrm{tr}_{E}\left(\nabla_{i}^{h_{0}}H\cdot H^{-1}\cdot\overline{{\partial}}_{i}H\right)=h_{0}\left(\nabla_{i}^{h_{0}}H\cdot H^{-1},\nabla^{h_{0}}_{i}H\right).$
Indeed, if we write $\nabla_{i}$ for $\nabla_{i}^{h_{0}}$, then for any $A\in\Gamma(M,E^{*}\otimes E)$, one has
$h_{0}(\nabla_{i}H\cdot A,\nabla_{i}H)=h_{0}^{\alpha\overline{p}}h_{0,\gamma\overline{q}}\left(\nabla_{i}H_{\alpha}^{\beta}\right)A_{\beta}^{\gamma}\left(\overline{\nabla_{i}H_{p}^{q}}\right),\quad\mathrm{tr}_{E}(\nabla_{i}H\cdot A\cdot\overline{{\partial}}_{i}H)=\left(\nabla_{i}H_{\alpha}^{\beta}\right)A_{\beta}^{\gamma}\left(\overline{{\partial}}_{i}H_{\gamma}^{\alpha}\right).$
We only need to show
(4.19) $\overline{{\partial}}_{i}H_{\gamma}^{\alpha}=h_{0}^{\alpha\overline{p}}h_{0,\gamma\overline{q}}\left(\overline{\nabla_{i}H_{p}^{q}}\right)$
Actually,
$\nabla_{i}H_{p}^{q}$ $=$ ${\partial}_{i}H_{p}^{q}-H_{t}^{q}h_{0}^{t\overline{\delta}}({\partial}_{i}h_{0,p\overline{\delta}})+H_{p}^{s}h_{0}^{q\overline{\varepsilon}}({\partial}_{i}h_{0,s\overline{\varepsilon}})$
$=$ ${\partial}_{i}H_{p}^{q}+H_{t}^{q}({\partial}_{i}h_{0}^{t\overline{\delta}})h_{0,p\overline{\delta}}+H_{p}^{s}h_{0}^{q\overline{\varepsilon}}({\partial}_{i}h_{0,s\overline{\varepsilon}}),$
and so
$h_{0}^{\alpha\overline{p}}h_{0,\gamma\overline{q}}\left(\overline{\nabla_{i}H_{p}^{q}}\right)$ $=$ $h_{0}^{\alpha\overline{p}}h_{0,\gamma\overline{q}}\left(\overline{{\partial}_{i}H_{p}^{q}}+\overline{H}{}_{t}^{q}(\overline{{\partial}}_{i}h_{0}^{\delta\overline{t}})h_{0,\delta\overline{p}}+\overline{H}{}_{p}^{s}h_{0}^{\varepsilon\overline{q}}(\overline{{\partial}}_{i}h_{0,\varepsilon\overline{s}})\right)$
$=$ $h_{0}^{\alpha\overline{p}}h_{0,\gamma\overline{q}}\overline{{\partial}_{i}H_{p}^{q}}+\overline{H}{}_{t}^{q}(\overline{{\partial}}_{i}h_{0}^{\alpha\overline{t}})h_{0,\gamma\overline{q}}+\overline{H}{}_{p}^{s}h_{0}^{\alpha\overline{p}}(\overline{{\partial}}_{i}h_{0,\gamma\overline{s}})$
$=$ $\overline{{\partial}}_{i}\left(h_{0}^{\alpha\overline{p}}h_{0,\gamma\overline{q}}\overline{H}{}_{p}^{q}\right)=\overline{{\partial}}_{i}H_{\gamma}^{\alpha},$
where the last identity holds since $H$ is Hermitian with respect to $h_{0}$.
Therefore, by (4.18) and (4.9), one obtains
$\mathrm{tr}_{E}\left(\Lambda_{\omega_{g}}\left(\sqrt{-1}{\partial}^{h_{0}}H\cdot H^{-1}\cdot{\overline{\partial}}H\right)\right)$ $=$ $\sum_{i}h_{0}\left(\nabla_{i}^{h_{0}}H\cdot H^{-1},\nabla^{h_{0}}_{i}H\right)$
$\geq$ $C_{4}^{-1}\sum_{i=1}^{n}|\nabla^{h_{0}}_{i}H\cdot H^{-1}|_{h}^{2}$
$=$ $C_{4}^{-1}|{\partial}^{h_{0}}H\cdot H^{-1}|_{h}^{2}.$
This completes the estimate (4.12). ∎
Proposition 4.4.
If there exists a constant $C_{4}\geq 1$ such that
(4.20) $C_{4}^{-1}h_{0}\leq h\leq C_{4}h_{0},$
then
(4.21) $\Delta_{g}\mathrm{tr}_{E}H\geq C_{4}^{-1}|{\partial}^{h_{0}}H\cdot H^{-1}|^{2}_{h}-\mathrm{tr}_{E}\Phi,$
where $\Delta_{g}=\mathrm{tr}_{\omega_{g}}\sqrt{-1}{\partial}{\overline{\partial}}$ on smooth functions.
Proof.
A straightforward computation gives
${\partial}^{*,h_{0}}({\partial}^{h_{0}}H\cdot H^{-1})\cdot H$ $=$ $\sqrt{-1}\Lambda_{\omega_{g}}\overline{{\partial}}({\partial}^{h_{0}}H\cdot H^{-1})\cdot H$
$=$ ${\partial}^{*,h_{0}}{\partial}^{h_{0}}H-\sqrt{-1}\Lambda_{\omega_{g}}({\partial}^{h_{0}}H\cdot\overline{{\partial}}H^{-1})\cdot H$
$=$ $\Delta_{{\partial}^{h_{0}}}H+\sqrt{-1}\Lambda_{\omega_{g}}({\partial}^{h_{0}}H\cdot H^{-1}\cdot\overline{{\partial}}H).$
On the other hand, by (3.16), one has
(4.23) $\Phi-\Omega\cdot H={\partial}^{*,h_{0}}({\partial}^{h_{0}}H\cdot H^{-1})\cdot H.$
Moreover,
$\mathrm{tr}_{E}H=h_{0}(H,\mathrm{Id}_{E})$
and so
$\Delta_{g}\mathrm{tr}_{E}H$ $=$ $\Lambda_{\omega_{g}}\sqrt{-1}{\partial}{\overline{\partial}}h_{0}(H,\mathrm{Id}_{E})$
$=$ $-\sqrt{-1}\Lambda_{\omega_{g}}h_{0}({\overline{\partial}}{\partial}^{h_{0}}H,\mathrm{Id}_{E})$
$=$ $-h_{0}\left(\Delta_{{\partial}^{h_{0}}}H,\mathrm{Id}_{E}\right)=-\mathrm{tr}_{E}\left(\Delta_{{\partial}^{h_{0}}}H\right).$
By taking the trace over $E$ of the equation in (4), one obtains
$\Delta_{\omega_{g}}\mathrm{tr}_{E}H=-\mathrm{tr}_{E}(\Delta_{{\partial}^{h_{0}}}H)$ $=$ $\mathrm{tr}_{E}\left(\sqrt{-1}\Lambda_{\omega_{g}}({\partial}^{h_{0}}H\cdot H^{-1}\cdot\overline{{\partial}}H)\right)-\mathrm{tr}_{E}({\partial}^{*,h_{0}}({\partial}^{h_{0}}H\cdot H^{-1})\cdot H)$
$\geq$ $C_{4}^{-1}|{\partial}^{h_{0}}H\cdot H^{-1}|^{2}_{h}+\mathrm{tr}_{E}(\Omega\cdot H-\Phi)$
$\geq$ $C_{4}^{-1}|{\partial}^{h_{0}}H\cdot H^{-1}|^{2}_{h}-\mathrm{tr}_{E}\Phi,$
where the first inequality follows from (4.12) and the second inequality holds since $\Omega>0$ with respect to $h_{0}$. ∎
Let ${\mathscr{E}}:=T^{*1,0}M\otimes E^{*}\otimes E$ be the Hermitian holomorphic vector bundle with the Hermitian metric induced by $g$ on $T^{*1,0}M$ and $h$ on $E$. Let $T$ be the tensor
(4.24) $T:={\partial}^{h_{0}}H\cdot H^{-1}\in\Omega^{1,0}(M,E^{*}\otimes E)\simeq\Gamma(M,{\mathscr{E}}).$
By using similar ideas as in the higher-order estimates for the Calabi-Yau equation, we derive the following estimate:
Proposition 4.5.
If there exists a constant $C_{4}\geq 1$ such that
(4.25) $C_{4}^{-1}h_{0}\leq h\leq C_{4}h_{0},$
then one has
(4.26) $\Delta_{\omega_{g}}|T|_{h}^{2}\geq-C_{10}\left(|T|^{2}_{h}+1\right)$
where $C_{10}$ depends on $M,\omega_{g},C_{4},h_{0},r$ and $\|\Phi\|_{C^{1}(M,\omega_{g},h_{0})}$.
Proof.
The Bochner-Kodaira formula establishes that
(4.27) $\Delta_{\omega_{g}}|T|_{h}^{2}=|{\partial}_{\mathscr{E}}T|_{h}^{2}+|\overline{{\partial}}T|_{h}^{2}+2\mathrm{Re}\left(\left\langle\mathrm{tr}_{\omega_{g}}\left(\sqrt{-1}{\partial}_{\mathscr{E}}\overline{{\partial}}_{\mathscr{E}}T\right),T\right\rangle_{\mathscr{E}}\right)-\mathrm{Ric}^{\mathscr{E}}(T,T),$
where $\mathrm{Ric}^{\mathscr{E}}$ is the Hermitian-Yang-Mills tensor of $\mathscr{E}$}. It is clear that
(4.28) $\mathrm{tr}_{\omega_{g}}\left(\sqrt{-1}{\partial}_{\mathscr{E}}\overline{{\partial}}_{\mathscr{E}}T\right)=\left(g^{i\overline{j}}\nabla^{h}_{i}\nabla^{h}_{\overline{j}}T_{k\alpha}^{\beta}\right)\;dz^{k}\otimes e^{\alpha}\otimes e_{\beta},$
where $\nabla^{h}$ is the Chern connection on $\mathscr{E}$}. Note that, we computed in (3.9) that
(4.29) $T=T_{k\alpha}^{\beta}\;dz^{k}\otimes e^{\alpha}\otimes e_{\beta}=\left((\Gamma^{h})_{i\alpha}^{\beta}-(\Gamma^{h_{0}})_{i\alpha}^{\beta}\right)dz^{i}\otimes e^{\alpha}\otimes e_{\beta}.$
Hence,
[…]
$=$ $g^{i\overline{j}}\nabla^{h}_{i}(R^{h})_{k\overline{j}\alpha}{}^{\beta}-g^{i\overline{j}}(\nabla_{i}^{h}-\nabla_{i}^{h_{0}})(R^{h_{0}})_{k\overline{j}\alpha}{}^{\beta}-g^{i\overline{j}}\nabla^{h_{0}}_{i}(R^{h_{0}})_{k\overline{j}\alpha}{}^{\beta}$,
where $\nabla^{h_{0}}$ is the Chern connection defined by $h_{0}$ and $\omega_{g}$. By using the Bianchi identity, one has
(4.31) $\nabla^{h}_{i}(R^{h})_{k\overline{j}\alpha}{}^{\beta}=\nabla^{h}_{k}(R^{h})_{i\overline{j}\alpha}{}^{\beta}$,
and so we get
$-g^{i\overline{j}}\nabla^{h}_{i}\nabla^{h}_{\overline{j}}T_{k\alpha}^{\beta}$ $=$ $g^{i\overline{j}}\nabla^{h}_{k}(R^{h})_{i\overline{j}\alpha}{}^{\beta}-g^{i\overline{j}}(\nabla_{i}^{h}-\nabla_{i}^{h_{0}})(R^{h_{0}})_{k\overline{j}\alpha}{}^{\beta}-g^{i\overline{j}}\nabla^{h_{0}}_{k}(R^{h_{0}})_{i\overline{j}\alpha}{}^{\beta}$
$=$ $\nabla^{h}_{k}(S^{h})_{\alpha}^{\beta}-g^{i\overline{j}}(\nabla_{i}^{h}-\nabla_{i}^{h_{0}})(R^{h_{0}})_{k\overline{j}\alpha}{}^{\beta}-\nabla^{h_{0}}_{k}(S^{h_{0}})_{\alpha}{}^{\beta}$,
By using the identity $S^{h}=\Phi\cdot H^{-1}$, we get
(4.32) $\nabla^{h}_{k}(S^{h})_{\alpha}^{\beta}=\nabla^{h}_{k}(\Phi\cdot H^{-1})_{\alpha}^{\beta}=\left(\nabla^{h}_{k}\Phi_{\alpha}^{\gamma}\right)(H^{-1})_{\gamma}^{\beta}+\Phi_{\alpha}^{\gamma}\nabla^{h}_{k}(H^{-1})_{\gamma}^{\beta}.$
Since $h_{0}$ and $h$ are equivalent,
(4.33) $|H^{-1}|_{h}\leq\sqrt{r}C_{4}$
and
(4.34) $\nabla^{h}_{k}\Phi_{\alpha}^{\gamma}=\nabla^{h_{0}}_{k}\Phi_{\alpha}^{\gamma}+(\nabla^{h}_{k}-\nabla^{h_{0}}_{k})\Phi_{\alpha}^{\gamma},$
we conclude that
(4.35) $|\left(\nabla^{h}_{k}\Phi_{\alpha}^{\gamma}\right)(H^{-1})_{\gamma}^{\beta}|_{h}\leq\widetilde{C}_{5}|{\partial}^{h_{0}}\Phi|_{h}+\widetilde{C}_{6}|T|_{h}|\Phi|_{h}.$
Similarly,
(4.36) $\nabla_{k}^{h}(H^{-1})^{\alpha}_{\gamma}=\nabla_{k}^{h}(h^{\alpha\overline{\beta}}h_{0,\gamma\overline{\beta}})=h^{\alpha\overline{\beta}}\nabla_{k}^{h}h_{0,\gamma\beta}=h^{\alpha\overline{\beta}}\left(\nabla_{k}^{h}-\nabla_{k}^{h_{0}}\right)h_{0,\gamma\overline{\beta}}.$
This implies
(4.37) $|\nabla^{h}_{k}(S^{h})_{\alpha}^{\beta}|_{h}\leq C_{5}|{\partial}^{h_{0}}\Phi|_{h}+C_{6}|T|_{h}|\Phi|_{h}$
where $\widetilde{C}_{5},\widetilde{C}_{6},$ $C_{5}$ and $C_{6}$ are constants depending on $M,\omega_{g},h_{0},r$ and $C_{4}$. It is easy to see that
(4.38) $|(\nabla_{i}^{h}-\nabla_{i}^{h_{0}})(R^{h_{0}})_{k\overline{j}\alpha}{}^{\beta}|_{h}\leq C_{7}|T|_{h}$
and
(4.39) $|\nabla^{h_{0}}_{k}(S^{h_{0}})_{\alpha}{}^{\beta}|_{h}\leq C_{8}$
where $C_{7}$ and $C_{8}$ are constants depending on $M,\omega_{g},h_{0}$ and $C_{4}$. By (4.28), (4.30), (4.37), (4.38) and (4.39), we deduce that
(4.40) $\left|\mathrm{tr}_{\omega_{g}}\left(\sqrt{-1}{\partial}_{\mathscr{E}}\overline{{\partial}}_{\mathscr{E}}T\right)\right|_{h}\leq C_{5}|{\partial}^{h_{0}}\Phi|_{h}+C_{6}|T|_{h}|\Phi|_{h}+C_{7}|T|_{h}+C_{8}.$
Note that ${\mathscr{E}}=T^{*1,0}M\otimes E^{*}\otimes E$ and the Hermitian-Yang-Mills tensor $\mathrm{Ric}^{\mathscr{E}}$ is a linear combination of the Hermitian-Yang-Mills tensors $\mathrm{Ric(\omega_{g})}$ and $S^{h}$, i.e.,
(4.41) $\mathrm{Ric}^{\mathscr{E}}=-\mathrm{Ric(\omega_{g})}\otimes\mathrm{Id}_{E^{*}\otimes E}+\mathrm{Id}_{T^{*1,0}M}\otimes\left(-\left(S^{h}\right)^{t}\otimes\mathrm{Id}_{E}+\mathrm{Id}_{E^{*}}\otimes S^{h}\right).$
Since $h$ and $h_{0}$ are equivalent and $\Phi=S^{h}\cdot H$, the following estimate holds
(4.42) $\mathrm{Ric}^{\mathscr{E}}(T,T)\leq C_{9}|T|^{2}_{h}\left(1+|\Phi|_{h}\right).$
By (4.27), (4.40) and (4.42), one has
$\Delta_{g}|T|_{h}^{2}\geq-\left(2C_{5}|{\partial}^{h_{0}}\Phi|_{h}|T|_{h}+2C_{6}|T|^{2}_{h}|\Phi|_{h}+2C_{7}|T|^{2}_{h}+2C_{8}|T|_{h}+C_{9}|T|^{2}_{h}+C_{9}|T|^{2}_{h}|\Phi|_{h}\right).$
By using the Cauchy-Schwarz inequality, we establish
(4.43) $\Delta_{\omega_{g}}|T|_{h}^{2}\geq-C_{10}\left(|T|^{2}_{h}+1\right)$
where $C_{10}$ depends on $M,\omega_{g},C_{4},h_{0},r$ and $\|\Phi\|_{C^{1}(M,\omega_{g},h_{0})}$. ∎
Proof of Theorem 4.1.
By Proposition 4.2, there exists $C_{4}=C_{4}(M,\omega_{g},h_{0},C_{1})\geq 1$ such that
(4.44) $C_{4}^{-1}h_{0}\leq h\leq C_{4}h_{0}.$
Moreover,
(4.45) $\mathrm{tr}_{E}\Phi\leq\sqrt{r}|\Phi|_{h_{0}}$
By Proposition 4.4 and Proposition 4.5,
(4.46) $\Delta_{g}\mathrm{tr}_{E}H\geq C_{4}^{-1}|T|^{2}_{h}-\mathrm{tr}_{E}\Phi\geq C_{4}^{-1}|T|^{2}_{h}-\sqrt{r}|\Phi|_{h_{0}},$
and
(4.47) $\Delta_{g}|T|_{h}^{2}\geq-C_{10}(|T|_{h}^{2}+1),$
For a large $L>0$ such that $LC_{4}^{-1}-C_{10}>1$, one has
(4.48) $\Delta_{g}(|T|_{h}^{2}+L\mathrm{tr}_{E}H)\geq(LC_{4}^{-1}-C_{10})|T|_{h}^{2}-(LC_{11}+C_{10}),$
where $C_{11}$ is an upper bound of $\sqrt{r}|\Phi|_{h_{0}}$. At a maximal point $p\in M$ of $(|T|_{h}^{2}+L\mathrm{tr}_{E}H)$, we obtain the estimate
(4.49) $|T|_{h}^{2}(p)\leq C_{12}.$
Since $\mathrm{tr}_{E}H\leq rC_{4}$, we get
(4.50) $(|T|_{h}^{2}+L\mathrm{tr}_{E}H)(p)\leq C_{13}.$
Therefore, for any point $x\in M$, one has
(4.51) $|T|^{2}_{h}(x)\leq(|T|_{h}^{2}+L\mathrm{tr}_{E}H)(x)\leq(|T|_{h}^{2}+L\mathrm{tr}_{E}H)(p)\leq C_{13}.$
Since $T={\partial}^{h_{0}}H\cdot H^{-1}$,
(4.52) $|{\partial}^{h_{0}}H\cdot H^{-1}|^{2}_{h}\leq C_{13}.$
By (4.10), there exists $C_{14}=C(M,\omega_{g},h_{0},C_{1},r,\|\Phi\|_{C^{1}(M,\omega_{g},h_{0})})>0$ such that
(4.53) $|{\partial}^{h_{0}}H|_{h_{0}}^{2}\leq C_{14}.$
Moreover, by the relation (4.19), we deduce that
(4.54) $|{\overline{\partial}}H|_{h_{0}}^{2}\leq C_{15}.$
In summary, we establish the uniform $C^{1}$-estimate $|H|_{C^{1}(M,\omega_{g},h_{0})}\leq C_{16}$ where $C_{16}$ depends on $M,\omega_{g},h_{0},C_{1},r$ and $\|\Phi\|_{C^{1}(M,\omega_{g},h_{0})}$. ∎
5. Proofs of main results
In this section we prove Theorem 1.1:
Theorem 5.1.
Let $E$ be a holomorphic vector bundle over a compact Kähler manifold $(M,\omega_{g})$. Suppose that there exists a smooth Hermitian metric $h_{0}$ on $E$ such that $\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}>0$. Then for any Hermitian positive definite tensor $P\in\Gamma(M,E^{*}\otimes\overline{E}^{*})$, there exists a unique smooth Hermitian metric $h$ on $E$ such that
(5.1) $\Lambda_{\omega_{g}}\sqrt{-1}R^{h}=P.$
Remark 5.2.
The dual equation case can be addressed via Serre duality. Suppose that there exists a smooth Hermitian metric $h_{0}$ on $E$ such that $\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}<0$. Then the dual metric $h_{0}^{*}$ on $E^{*}$ satisfies
$\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}^{*}}>0$
as a Hermitian tensor in $\Gamma(M,E\otimes\overline{E})$ under the natural identification $E\cong(E^{*})^{*}$. Hence, for any Hermitian positive definite tensor $P_{*}\in\Gamma(M,E\otimes\overline{E})$, there exists a unique smooth Hermitian metric $h_{*}$ on $E^{*}$ such that
(5.2) $\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{*}}=P_{*}.$
Or equivalently, for any Hermitian negative definite tensor $P\in\Gamma(M,E^{*}\otimes\overline{E}^{*})$, there exists a unqiue smooth Hermitian metric $h$ on $E$ satisfying the dual equation:
(5.3) $\Lambda_{\omega_{g}}\sqrt{-1}R^{h}=H\cdot P\cdot\overline{H^{t}}$
where $H=h\cdot h_{0}^{-1}\in\Gamma(M,E^{*}\otimes E)$.
We begin by proving the following result, which ensures the closedness property required in Theorem 5.1.
Theorem 5.3.
Let $(M,\omega_{g})$ be a compact Kähler manifold and $(E,h_{0})$ be a Hermitian holomorphic vector bundle with $S^{h_{0}}>0$. Suppose that $\{h_{m}\}$ is a sequence of smooth Hermitian metrics on $E$ and $S^{h_{m}}$ are the corresponding Hermitian-Yang-Mills tensors in $\Gamma(M,E^{*}\otimes\overline{E}^{*})$. If
(5.4) $\lim_{m}\|S^{h_{m}}-P\|_{C^{\infty}(M,\omega_{g},h_{0})}=0$
for some Hermitian positive definite $P\in\Gamma(M,E^{*}\otimes\overline{E}^{*})$, then there exists a unique smooth Hermitian metric $h$ on $E$ such that
(5.5) $\Lambda_{\omega_{g}}\sqrt{-1}R^{h}=P.$
Proof.
We fix several notations in $\Gamma(M,E^{*}\otimes E)$:
(5.6) $H_{m}=h_{0}^{\alpha\overline{\gamma}}h_{m,\beta\overline{\gamma}}\;e^{\beta}\otimes e_{\alpha},\quad\Phi_{m}=h_{0}^{\alpha\overline{\gamma}}(S^{h_{m}})_{\beta\overline{\gamma}}\;e^{\beta}\otimes e_{\alpha},\quad\Phi:=h_{0}^{\alpha\overline{\gamma}}P_{\beta\overline{\gamma}}\;e^{\beta}\otimes e_{\alpha}.$
Hence, $\Phi>0$ with respect to $h_{0}$. Since $\lim_{m}\|S^{h_{m}}-P\|_{C^{\infty}(M,\omega_{g},h_{0})}=0$, one has
(5.7) $\lim_{m}\|\Phi_{m}-\Phi\|_{C^{\infty}(M,\omega_{g},h_{0})}=0.$
In particular, there exist uniform constants $c_{1},c_{2}>0$ and $m_{0}>0$, depending on $M,\omega_{g},h_{0}$ and $P$, such that for any $m\geq m_{0}$, the following holds
(5.8) $c_{1}^{-1}\mathrm{Id}_{E}\leq\Phi_{m}\leq c_{1}\mathrm{Id}_{E},\quad\left\|\Phi_{m}\right\|_{C^{1}(M,\omega_{g},h_{0})}\leq c_{2}.$
By Theorem 4.1, there exists $C_{0}=C_{0}(M,\omega_{g},h_{0},r,c_{1},c_{2})>0$ such that,
(5.9) $\left\|H_{m}\right\|_{C^{1}(M,\omega_{g},h_{0})}\leq C_{0}.$
By the curvature relation (3.16), one deduces that
(5.10) $\Phi_{m}=\Omega\cdot H_{m}+{\partial}^{*,h_{0}}({\partial}^{h_{0}}H_{m}\cdot H_{m}^{-1})\cdot H_{m}.$
Moreover, by using similar arguments as in the proof of Proposition 4.4, one has
(5.11) $\Delta_{{\partial}^{h_{0}}}H_{m}=\Phi_{m}-\Omega\cdot H_{m}-\sqrt{-1}\Lambda_{\omega_{g}}({\partial}^{h_{0}}H_{m}\cdot H_{m}^{-1}\cdot\overline{{\partial}}H_{m}).$
Let
(5.12) $W_{m}:=\Phi_{m}-\Omega\cdot H_{m}-\sqrt{-1}\Lambda_{\omega_{g}}({\partial}^{h_{0}}H_{m}\cdot H_{m}^{-1}\cdot\overline{{\partial}}H_{m}).$
Hence,
(5.13) $\Delta_{{\partial}^{h_{0}}}H_{m}=W_{m}.$
By (5.8) and (5.9),
(5.14) $\|W_{m}\|_{C^{0}(M,\omega_{g},h_{0})}\leq C_{1}(M,\omega_{g},h_{0},P)$
For a fixed large integer $p$, there exists a constant $C_{2}=C_{2}(M,\omega_{g},h_{0},P,p)>0$ such that
(5.15) $\left\|W_{m}\right\|_{L^{p}(M,\omega_{g},h_{0})}\leq C_{2}.$
By $W^{2,p}$-estimate for elliptic equations, there exists $C_{3}=C_{3}(M,\omega_{g},h_{0},p)>0$ such that
(5.16) $\left\|H_{m}\right\|_{W^{2,p}(M,\omega_{g},h_{0})}\leq C_{3}\left(\left\|H_{m}\right\|_{L^{p}(M,\omega_{g},h_{0})}+\left\|W_{m}\right\|_{L^{p}(M,\omega_{g},h_{0})}\right).$
In particular, $\{H_{m}\}$ is uniformly bounded in $W^{2,p}(M,E^{*}\otimes E)$ with respect to the norm induced by $\omega_{g},h_{0}$. Since $p$ is large enough and $M$ is compact, by the compact embedding theorem, there exist a subsequence $\{H_{m_{i}}\}$, $\alpha\in(0,1)$ and $H\in C^{1,\alpha}(M,E^{*}\otimes E)$ such that
(5.17) $\lim_{i}\|H_{m_{i}}-H\|_{C^{1,\alpha}(M,E^{*}\otimes E)}=0.$
Moreover, $H$ is $h_{0}$-Hermitian positive definite. In particular, by using (5.12), one has
(5.18) $\lim_{i}\|W_{m_{i}}-W\|_{C^{0,\alpha}(M,E^{*}\otimes E)}=0$
where $W\in C^{0,\alpha}(M,E^{*}\otimes E)$ is given by
(5.19) $W=\Phi-\Omega\cdot H-\sqrt{-1}\Lambda_{\omega_{g}}({\partial}^{h_{0}}H\cdot H^{-1}\cdot\overline{{\partial}}H).$
By Schauder’s estimate, there exists $C_{4}=C_{4}(M,\omega_{g},h_{0},p,\alpha)>0$ such that
(5.20) $\left\|H_{m_{i}}-H_{m_{j}}\right\|_{C^{2,\alpha}(M,E^{*}\otimes E)}\leq C_{4}\left(\left\|H_{m_{i}}-H_{m_{j}}\right\|_{C^{0,\alpha}(M,E^{*}\otimes E)}+\left\|W_{m_{i}}-W_{m_{j}}\right\|_{C^{0,\alpha}(M,E^{*}\otimes E)}\right).$
In particular, $\{H_{m_{i}}\}$ is a Cauchy sequence in $C^{2,\alpha}(M,E^{*}\otimes E)$. Moreover, by (5.17), we deduce that $H\in C^{2,\alpha}(M,E^{*}\otimes E)$ and
(5.21) $\lim_{i}\|H_{m_{i}}-H\|_{C^{2,\alpha}(M,E^{*}\otimes E)}=0.$
When $m_{i}\rightarrow+\infty$ in (5.11), one has the following equation
(5.22) $\Delta_{{\partial}^{h_{0}}}H=W=\Phi-\Omega\cdot H-\sqrt{-1}\Lambda_{\omega_{g}}({\partial}^{h_{0}}H\cdot H^{-1}\cdot\overline{{\partial}}H).$
By the standard bootstrap process, one concludes that $H\in\Gamma(M,E^{*}\otimes E)$. When $m_{i}\rightarrow+\infty$ in (5.10),
(5.23) $\Phi=\Omega\cdot H+{\partial}^{*,h_{0}}({\partial}^{h_{0}}H\cdot H^{-1})\cdot H.$
If we set $h=H\cdot h_{0}$, then by (3.16), we obtain the Hermitian-Yang-Mills tensor relation:
(5.24) $h_{0}^{\gamma\overline{\beta}}S^{h}_{\alpha\overline{\beta}}\;e^{\alpha}\otimes e_{\gamma}=\Omega\cdot H+{\partial}^{*}({\partial}^{h_{0}}H\cdot H^{-1})\cdot H.$
By (5.23) and (5.24), we conclude that
(5.25) $S^{h}_{\alpha\overline{\beta}}=P_{\alpha\overline{\beta}}.$
This completes the proof. ∎
Before proving the openness part of Theorem 5.1, we fix some key notions. Let $E$ be a holomorphic vector bundle over a compact Kähler manifold $(M,\omega_{g})$. The subspace of smooth Hermitian sections in $\Gamma(M,E^{*}\otimes\overline{E}^{*})$ is denoted by $\mathrm{Herm}(E)$, and the subspace of smooth Hermitian metrics in $\Gamma(M,E^{*}\otimes\overline{E}^{*})$ is denoted by $\mathrm{Herm}^{+}(E)$. There is a natural map $G:\mathrm{Herm}^{+}(E)\rightarrow\mathrm{Herm}(E)$ given by the Hermitian-Yang-Mills tensor:
(5.26) $G(h)=S^{h}=S^{h}_{\alpha\overline{\beta}}e^{\alpha}\otimes\overline{e}^{\beta}\in\Gamma(M,E^{*}\otimes\overline{E}^{*}),$
where $S^{h}_{\alpha\overline{\beta}}=g^{i\overline{j}}R^{h}_{i\overline{j}\alpha\overline{\beta}}$. Fix a smooth Hermitian metric $h_{0}$ on $E$, we define
(5.27) $\mathrm{Herm}(E,h_{0}):=\left\{S\in\Gamma(M,E^{*}\otimes E)|\ S\text{ is $h_{0}$-Hermitian}\right\},$
and
(5.28) $\mathrm{Herm}^{+}(E,h_{0}):=\left\{S\in\mathrm{Herm}(E,h_{0})|\ S>0\right\}.$
It is clear that the map $F:\mathrm{Herm}(E)\rightarrow\mathrm{Herm}(E,h_{0})$ given by
(5.29) $F(h)=h\cdot h_{0}^{-1}$
is an isomorphism and its restriction $F:\mathrm{Herm}^{+}(E)\rightarrow\mathrm{Herm}^{+}(E,h_{0})$ is also an isomorphism. There is an induced map $\widetilde{G}:\mathrm{Herm}^{+}(E,h_{0})\rightarrow\mathrm{Herm}(E,h_{0})$
(5.30) $\widetilde{G}=F\circ G\circ F^{-1}.$
Since $\mathrm{Herm}(E,h_{0})$ is a closed linear subspace of $\Gamma(M,E^{*}\otimes E)$, for any $S\in\mathrm{Herm}(E,h_{0})$, there is an isomorphism
(5.31) $\mathrm{Herm}(E,h_{0})\simeq T_{S}\mathrm{Herm}(E,h_{0})$
mapping $\Psi\in\mathrm{Herm}(E,h_{0})$ to $\gamma_{\Psi}]\in T_{S}\mathrm{Herm}(E,h_{0})$, where $\gamma_{\Psi}$ is a curve $\gamma_{\Psi}(t)=S+t\Psi$ in $\mathrm{Herm}^{+}(E,h_{0})$.
Lemma 5.4.
For any $H\in\mathrm{Herm}^{+}(E,h_{0})$,
(5.32) $\widetilde{G}(H)={\partial}^{*,h_{0}}({\partial}^{h_{0}}H\cdot H^{-1})\cdot H+\Omega\cdot H,$
where $\Omega=S^{h_{0}}\in\Gamma(M,E^{*}\otimes E)$. Moreover, the linearization map
(5.33) $\mathscr{L}_{\mathrm{Id}_{E}}:\mathrm{Herm}(E,h_{0})\rightarrow\mathrm{Herm}(E,h_{0})$
of $\widetilde{G}$ at the point $\mathrm{Id}_{E}$ is given by
(5.34) $\mathscr{L}_{\mathrm{Id}_{E}}(\Psi)=\Delta_{{\partial}^{h_{0}}}\Psi+\Omega\cdot\Psi.$
Proof.
For any $H\in\mathrm{Herm}^{+}(E,h_{0})$, it induces a Hermitian metric
(5.35) $h=F^{-1}(H)=H\cdot h_{0}\in\mathrm{Herm}^{+}(E).$
By (3.6), one has the difference of the Hermitian-Yang-Mills tensors:
(5.36) $(S^{h})_{\alpha}^{\beta}-(S^{h_{0}})_{\alpha}^{\beta}=({\partial}^{*,h_{0}}({\partial}^{h_{0}}H\cdot H^{-1}))_{\alpha}^{\beta}.$
Hence,
(5.37) $S^{h}_{\alpha\overline{\gamma}}=h_{\beta\overline{\gamma}}(S^{h_{0}})_{\alpha}^{\beta}+h_{\beta\overline{\gamma}}({\partial}^{*,h_{0}}({\partial}^{h_{0}}H\cdot H^{-1}))_{\alpha}^{\beta}.$
That is
(5.38) $G(F^{-1}(H))=S^{h}_{\alpha\overline{\gamma}}\;e^{\alpha}\otimes\overline{e}{}^{\gamma}=\left(h_{\beta\overline{\gamma}}(S^{h_{0}})_{\alpha}^{\beta}+h_{\beta\overline{\gamma}}({\partial}^{*,h_{0}}({\partial}^{h_{0}}H\cdot H^{-1}))_{\alpha}^{\beta}\right)\;e^{\alpha}\otimes\overline{e}{}^{\gamma}$.
Therefore, one has
$\widetilde{G}(H)=F(G(F^{-1}(H)))$ $=$ $h_{0}^{\delta\overline{\gamma}}h_{\beta\overline{\gamma}}\left((S^{h_{0}})_{\alpha}^{\beta}+({\partial}^{*}({\partial}^{h_{0}}H\cdot H^{-1}))_{\alpha}^{\beta}\right)\;e^{\alpha}\otimes e_{\delta}$
$=$ ${\partial}^{*,h_{0}}({\partial}^{h_{0}}H\cdot H^{-1})\cdot H+\Omega\cdot H.$
We compute the tangent map of $\widetilde{G}$ at $\mathrm{Id}_{E}\in\mathrm{Herm}^{+}(E,h_{0})$. For any $\Psi\in\mathrm{Herm}(E,h_{0})$, there is a curve $\gamma(t)=\mathrm{Id}_{E}+t\Psi$ in $\mathrm{Herm}^{+}(E,h_{0})$ when $t$ is small. It is easy to see that
(5.39) $(\mathrm{Id}_{E}+t\Psi)^{-1}=\mathrm{Id}_{E}-t\Psi+O(t^{2}).$
Since ${\partial}^{h_{0}}(\mathrm{Id}_{E})=0$, one has
$\widetilde{G}(\gamma(t))$ $=$ $t{\partial}^{*,h_{0}}(({\partial}^{h_{0}}\Psi)\cdot(\mathrm{Id}_{E}-t\Psi))+\Omega+t\Omega\cdot\Psi+O(t^{2})$
$=$ $\Omega+t\left(\Delta_{{\partial}^{h_{0}}}\Psi+\Omega\cdot\Psi\right)+O(t^{2}).$
Hence, the linearization of $\widetilde{G}$ at $\mathrm{Id}_{E}$ is
(5.40) $\mathscr{L}_{\mathrm{Id}_{E}}(\Psi)=\Delta_{{\partial}^{h_{0}}}\Psi+\Omega\cdot\Psi.$
This completes the proof. ∎
The following result is useful in the $C^{k,\alpha}$ and $W^{k,p}$ estimates:
Lemma 5.5.
Let $S\in\mathrm{Herm}(E,h_{0})$ and $\Omega\in\mathrm{Herm}^{+}(E,h_{0})$. If $\Psi\in\Gamma(M,E^{*}\otimes E)$ satisfies
(5.41) $\Delta_{{\partial}^{h_{0}}}\Psi+\Omega\cdot\Psi=S,$
then $\Psi\in\mathrm{Herm}(E,h_{0})$.
Proof.
Let $T\in\Gamma(M,E^{*}\otimes E)$. We show that
(5.42) $\left({\partial}^{h_{0}}T\right)^{*}={\overline{\partial}}T^{*},\quad\left({\overline{\partial}}T\right)^{*}={\partial}^{h_{0}}T^{*}.$
Here and in the following, the Hermitian adjoint is taken with respect to $h_{0}$. Indeed,
(5.43) ${\partial}\left(h_{0}(Tv,w)\right)=h_{0}({\partial}^{h_{0}}T\cdot v+T\cdot{\partial}^{h_{0}}v,w)+h_{0}(Tv,{\overline{\partial}}w),$
for any $v,w\in\Gamma(M,E)$. Similarly, one has
(5.44) ${\partial}\left(h_{0}(v,T^{*}w)\right)=h_{0}({\partial}^{h_{0}}v,T^{*}w)+h_{0}(v,{\overline{\partial}}T^{*}\cdot w+T^{*}{\overline{\partial}}w).$
Since $h_{0}(Tv,w)=h_{0}(v,T^{*}w)$, we deduce that
$h_{0}({\partial}^{h_{0}}T\cdot v,w)=h_{0}(v,{\overline{\partial}}T^{*}\cdot w)$
which implies $\left({\partial}^{h_{0}}T\right)^{*}={\overline{\partial}}T^{*}$. By using the expressions of
${\partial}{\overline{\partial}}(h_{0}(Tv,w))={\partial}{\overline{\partial}}(h_{0}(v,T^{*}w)),$
we get
(5.45) $\left({\partial}^{h_{0}}{\overline{\partial}}T\right)^{*}={\overline{\partial}}{\partial}^{h_{0}}T^{*},\quad\left({\overline{\partial}}{\partial}^{h_{0}}T\right)^{*}={\partial}^{h_{0}}{\overline{\partial}}T^{*}.$
In particular,
(5.46) $(\Delta_{\overline{\partial}}T)^{*}=\Delta_{{\partial}^{h_{0}}}T^{*},\quad\left(\Delta_{{\partial}^{h_{0}}}T\right)^{*}=\Delta_{{\overline{\partial}}}T^{*}.$
Since $\Omega$ is $h_{0}$-Hermitian,
(5.47) $(\Omega\cdot\Psi)^{*}=\Psi^{*}\cdot\Omega^{*}=\Psi^{*}\cdot\Omega.$
By the Bochner-Kodaira identity on vector bundle $E^{*}\otimes E$, one can see
(5.48) $\Delta_{\overline{{\partial}}}\Psi^{*}-\Delta_{{\partial}^{h_{0}}}\Psi^{*}=\Omega\cdot\Psi^{*}-\Psi^{*}\cdot\Omega.$
Therefore, by (5.46), (5.47) and (5.48), we obtain that
(5.49) $\Delta_{{\partial}^{h_{0}}}\Psi^{*}+\Omega\cdot\Psi^{*}=\Delta_{{\overline{\partial}}}\Psi^{*}+\Psi^{*}\cdot\Omega=(\Delta_{{\partial}^{h_{0}}}\Omega+\Omega\cdot\Psi)^{*}=S^{*}=S.$
From (5.41) and (5.49), one has
(5.50) $\Delta_{{\partial}^{h_{0}}}\left(\Psi^{*}-\Psi\right)+\Omega\cdot(\Psi^{*}-\Psi)=0.$
Since $\Omega$ is $h_{0}$-Hermitian positive definite, by pairing with $(\Psi^{*}-\Psi)$ in (5.50), one deduces that $\Psi^{*}-\Psi=0$ and so $\Psi$ is $h_{0}$-Hermitian. ∎
Theorem 5.6.
Let $P\in\mathrm{Herm}^{+}(E)$. If there exists a Hermitian metric $h_{1}\in\mathrm{Herm}^{+}(E)$ such that
(5.51) $P=G(h_{1})\in\Gamma(M,E^{*}\otimes\overline{E}^{*}),$
then there exist open neighborhoods $U$ of $P\in\mathrm{Herm}(E)$ and $V$ of $h_{1}\in\mathrm{Herm}^{+}(E)$ such that $G:U\rightarrow V$ is an isomorphism.
Proof.
We fix the center point $h_{1}\in\mathrm{Herm}^{+}(E)$ and so $F(h_{1})=\mathrm{Id}_{E}$. Since $\widetilde{G}=F\circ G\circ F^{-1}$,
(5.52) $d\widetilde{G}_{\mathrm{Id}_{E}}=(dF)_{G(h_{1})}\circ(dG)_{h_{1}}\circ(dF^{-1})_{F(h_{1})}$
Moreover, since $F$ is an isomorphism, we only need to verify that
$\mathscr{L}_{\mathrm{Id}_{E}}=d\widetilde{G}_{\mathrm{Id}_{E}}:\mathrm{Herm}(E,h_{1})\rightarrow\mathrm{Herm}(E,h_{1})$
is an isomorphism. By Lemma 5.4, for any $\Psi\in\mathrm{Herm}(E,h_{1})$,
(5.53) $\mathscr{L}_{\mathrm{Id}_{E}}(\Psi)=\Delta_{{\partial}^{h_{1}}}\Psi+\Omega_{1}\cdot\Psi.$
where $\Omega_{1}=P\cdot h_{1}^{-1}\in\Gamma(M,E^{*}\otimes E)$. Since both $\Omega_{1}$ and $\Psi$ are $h_{1}$-Hermitian, $\mathscr{L}_{\mathrm{Id}_{E}}$ is self-adjoint with respect to $\omega_{g}$ and $h_{1}$. Since $\Omega_{1}>0$, one can see clearly that $\mathscr{L}_{\mathrm{Id}_{E}}$ is injective. Moreover, since $M$ is compact, there exist $m_{1},m_{2}>0$ such that
(5.54) $m_{1}\cdot\mathrm{Id}_{E}\leq\Omega_{1}\leq m_{2}\cdot\mathrm{Id}_{E}$
with respect to $h_{1}$. Hence,
(5.55) $(\mathscr{L}_{\mathrm{Id}_{E}}\Psi,\Psi)_{h_{1}}\geq(\Omega_{1}\cdot\Psi,\Psi)_{h_{1}}\geq m_{1}(\Psi,\Psi)_{h_{1}}.$
By applying the Lax-Milgram Theorem framework together with standard elliptic regularity theory, we conclude that $\mathscr{L}_{\mathrm{Id}_{E}}$ is an isomorphism. For the convenience of readers, we outline key details of the proof below. For any positive integer $m$, let
(5.56) $W^{m,2}(M,E^{*}\otimes E)$
be the space of $W^{m,2}$ sections of $E^{*}\otimes E$ with respect to the metric $\omega_{g}$ and $h_{1}$. Consider the pairing
(5.57) $\mathscr{B}:W^{1,2}(M,E^{*}\otimes E)\times W^{1,2}(M,E^{*}\otimes E)\rightarrow\mathbb{C}$
defined by
(5.58) $\mathscr{B}(\Psi,T)=(\mathscr{L}_{\mathrm{Id}_{E}}(\Psi),T)_{h_{1}}$
in the weak sense. It is clear that
(5.59) $\mathscr{B}(\Psi,T)=({\partial}^{h_{1}}\Psi,{\partial}^{h_{1}}T)+(\Omega_{1}\cdot\Psi,T)$
and so
(5.60) $|\mathscr{B}(\Psi,T)|\leq C_{1}\left\|\Psi\right\|_{W^{1,2}}\left\|T\right\|_{W^{1,2}},\quad|\mathscr{B}(\Psi,\Psi)|\geq C_{2}\left\|\Psi\right\|_{W^{1,2}}^{2},$
where $C_{1}=C_{1}(M,\omega_{g},h_{1})>0$ and $C_{2}=C_{2}(M,\omega_{g},h_{1})>0$. Since $\mathscr{L}_{\mathrm{Id}_{E}}$ is self-adjoint, by the Lax-Milgram Theorem, one establishes an isomorphism
(5.61) $\mathscr{L}_{\mathrm{Id}_{E}}:W^{1,2}(M,E^{*}\otimes E)\simeq\left(W^{1,2}(M,E^{*}\otimes E)\right)^{*}.$
We claim that for all $k\geq 1$,
(5.62) $\mathscr{L}_{\mathrm{Id}_{E}}:W^{k+2,2}(M,E^{*}\otimes E)\rightarrow W^{k,2}(M,E^{*}\otimes E)$
are isomorphisms. We only need to show the injectivity and surjectivity.
  1. (1)
    (Injectivity) Suppose that there exists some $\Psi\in W^{k+2,2}(M,E^{*}\otimes E)$ such that
    (5.63) $\mathscr{L}_{\mathrm{Id}_{E}}(\Psi)=0\in W^{k,2}(M,E^{*}\otimes E),$
    then by pairing with $\Psi$, one deduces $\Psi=0$.
  2. (2)
    (Surjectivity) There is a natural inclusion (e.g. [Bre11, p. 291])
    (5.64) $W^{k,2}(M,E\otimes E^{*})\subset W^{1,2}(M,E\otimes E^{*})\subset\left(W^{1,2}(M,E\otimes E^{*})\right)^{*}.$
    By (5.61), for each $S\in W^{k,2}(M,E^{*}\otimes E)$, there exists some
    (5.65) $\Psi\in W^{1,2}(M,E^{*}\otimes E)$
    such that
    (5.66) $\mathscr{L}_{\mathrm{Id}_{E}}(\Psi)=\Delta_{{\partial}^{h_{1}}}\Psi+\Omega_{1}\cdot\Psi=S$
    in the weak $W^{1,2}(M,E^{*}\otimes E)$ sense. By standard regularity theory for elliptic PDEs (e.g. [Eva10, p. 329]), one concludes that $\Psi\in W^{k+2,2}(M,E^{*}\otimes E)$. Hence, $\mathscr{L}_{\mathrm{Id}_{E}}$ in (5.62) is surjective.
By the Sobolev embedding theorem and Lemma 5.5, we deduce that the linearization map $\mathscr{L}_{\mathrm{Id}_{E}}:\mathrm{Herm}(E,h_{1})\rightarrow\mathrm{Herm}(E,h_{1})$ is an isomorphism and so Theorem 5.6 holds by the implicit function theorem. ∎
By refining the openness and closedness arguments in the proofs of Theorem 5.3 and Theorem 5.6, we establish an existence result under the $C^{k,\alpha}$-norm. For any $k>1$ and $\alpha\in(0,1)$, let $C^{k,\alpha}\mathrm{Herm}(E)$ and $C^{k,\alpha}\mathrm{Herm}^{+}(E)$ be $C^{k,\alpha}$ Hermitian sections and $C^{k,\alpha}$ Hermitian metrics on $E$ respectively. Here, the $C^{k,\alpha}$-norm is taken with respect to $\omega_{g}$. For any $h_{1}\in C^{k,\alpha}\mathrm{Herm}^{+}(E)$, one can also define $C^{\ell,\alpha}\mathrm{Herm}(E,h_{1})$ and $C^{\ell,\alpha}\mathrm{Herm}^{+}(E,h_{1})$ for any non-negative integer $\ell\leq k$. Moreover, we have the following diagram, which refines the construction in the proof of Theorem 5.6:
[…]
Theorem 5.7.
Let $E$ be a holomorphic vector bundle over a compact Kähler manifold $(M,\omega_{g})$. Suppose that there exists a smooth Hermitian metric $h_{0}$ on $E$ such that the Hermitian-Yang-Mills tensor $S^{h_{0}}>0$. Then for any integer $k>1$, $\alpha\in(0,1)$ and any $P\in C^{k,\alpha}{\mathrm{Herm}^{+}(E)}$, there exists a unique Hermitian metric $h\in C^{k+2,\alpha}{\mathrm{Herm}^{+}(E)}$ such that its Hermitian-Yang-Mills tensor $S^{h}$ coincides with $P$.
Proof.
The uniqueness of metric $h$ is established in Corollary 3.3 since $h$ is $C^{3}$. For the existence of such a metric, we define
(5.67) $\mathscr{R}=\{P\in C^{k,\alpha}{\mathrm{Herm}^{+}(E)}|\ S^{h}=P\text{ for some}\ h\in C^{k+2,\alpha}{\mathrm{Herm}^{+}(E)}\}.$
By a similar argument as in the proof of Theorem 5.6, one deduces that $\mathscr{R}$ is open in $C^{k,\alpha}{\mathrm{Herm}^{+}(E)}$. Actually, for any $h_{1}\in C^{k+2,\alpha}{\mathrm{Herm}^{+}(E)}$ with $S^{h_{1}}>0$, one can show that the linearization map
(5.68) $\mathscr{L}_{\mathrm{Id}_{E}}:C^{k+2,\alpha}{\mathrm{Herm}(E,h_{1})}\rightarrow C^{k,\alpha}{\mathrm{Herm}(E,h_{1})}$
of $\widetilde{G}_{k,\alpha}$ at point $\mathrm{Id}_{E}$ is given by
(5.69) $\mathscr{L}_{\mathrm{Id}_{E}}(\Psi)=\Delta_{{\partial}^{h_{1}}}\Psi+\Omega_{1}\cdot\Psi.$
Moreover, it is an isomorphism. Indeed, there is an inclusion
(5.70) $C^{k,\alpha}{\mathrm{Herm}(E,h_{1})}\subset W^{1,2}(M,E^{*}\otimes E)\subset\left(W^{1,2}(M,E^{*}\otimes E)\right)^{*}.$
By using similar argument as in the proof of Theorem 5.6, for any $S\in C^{k,\alpha}{\mathrm{Herm}(E,h_{1})}$, by (5.70), there exists some
(5.71) $\Psi\in W^{1,2}(M,E^{*}\otimes E)$
such that
(5.72) $\Delta_{{\partial}^{h_{1}}}\Psi+\Omega_{1}\cdot\Psi=S$
in the weak $W^{1,2}(M,E^{*}\otimes E)$ sense. By standard regularity theory for elliptic PDEs,
(5.73) $\Psi\in W^{k+2,2}(M,E^{*}\otimes E).$
By the Sobolev embedding theorem, $\Psi\in W^{3,p_{1}}(M,E^{*}\otimes E)$ with
(5.74) $\frac{1}{p_{1}}=\frac{1}{2}-\frac{k-1}{2n}$
By repeating the above process, we know
(5.75) $\Psi\in W^{k+2,p_{1}}(M,E^{*}\otimes E).$
By induction, one can show that
(5.76) $\Psi\in W^{k+2,p_{m}}(M,E^{*}\otimes E)$
where
(5.77) $\frac{1}{p_{m}}=\frac{1}{p_{m-1}}-\frac{k-1}{2n}.$
By the induction formulation,
(5.78) $\frac{1}{p_{m}}=\frac{1}{2}-\frac{m(k-1)}{2n}.$
Let $m_{0}=\lceil\frac{n}{k-1}\rceil-1$. Then
(5.79) $\frac{1}{p_{m_{0}}}\leq\frac{k-1}{2n}\leq\frac{k-\alpha}{2n}.$
In particular, by the Sobolev embedding theorem, we obtain
(5.80) $W^{k+2,p_{m_{0}}}(M,E^{*}\otimes E)\subset W^{k+2,\frac{2n}{k-\alpha}}(M,E^{*}\otimes E)\subset C^{2,\alpha}(M,E^{*}\otimes E).$
Hence, $\Psi\in C^{2,\alpha}(M,E^{*}\otimes E)$. Since $S\in C^{k,\alpha}(M,E^{*}\otimes E)$ and $\Omega_{1}\in C^{k,\alpha}(M,E^{*}\otimes E)$, by Schauder’s theory for the elliptic equation
$\Delta_{{\partial}^{h_{1}}}\Psi+\Omega_{1}\cdot\Psi=S,$
we conclude that $\Psi\in C^{k+2,\alpha}(M,E^{*}\otimes E)$. Hence by Lemma 5.5, the map in (5.68) is an isomorphism and so $\mathscr{R}$ is open in $C^{k,\alpha}{\mathrm{Herm}^{+}(E)}$. By using similar arguments as in the proof of Theorem 5.3, one can show that: if $\{h_{m}\}$ is a sequence of $C^{k+2,\alpha}$-Hermitian metrics on $E$ and
(5.81) $\lim_{m}\|S^{h_{m}}-P\|_{C^{k,\alpha}(M,\omega_{g},h_{0})}=0$
for some $P\in C^{k,\alpha}\mathrm{Herm}^{+}(E)$, then there exists a unique $h\in C^{k+2,\alpha}{\mathrm{Herm}^{+}(E)}$ such that $S^{h}=P$. Hence, $\mathscr{R}$ is closed in $C^{k,\alpha}{\mathrm{Herm}^{+}(E)}$. Since $C^{k,\alpha}{\mathrm{Herm}^{+}(E)}$ is connected, one deduces that $\mathscr{R}=C^{k,\alpha}{\mathrm{Herm}^{+}(E)}$. ∎
Proof of Theorem 5.1.
For any $P\in\mathrm{Herm}^{+}(E)$, by Theorem 5.7, there exists a unique Hermitian metric $h_{k}\in C^{k+2,\alpha}{\mathrm{Herm}^{+}(E)}$ such that
(5.82) $S^{h_{k}}=P.$
Moreover, by Corollary 3.3, $h_{k}$ are the same for all $k>1$ and we denote it by $h$. Hence, $h$ is a smooth Hermitian metric on $E$ and $S^{h}=P$. ∎
We demonstrate that Theorem 1.1 holds in a slightly more general setting.
Theorem 5.8.
Let $E$ be a holomorphic vector bundle over a compact Kähler manifold $(M,\omega_{g})$. Suppose that there exists a smooth Hermitian metric $h_{0}$ on $E$ such that the Hermitian-Yang-Mills tensor $\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}$ is quasi-positive definite. Then for any Hermitian positive definite tensor $P\in\Gamma(M,E^{*}\otimes\overline{E}^{*})$, there exists a unique smooth Hermitian metric $h$ on $E$ such that
(5.83) $\Lambda_{\omega_{g}}\sqrt{-1}R^{h}=P.$
Recall that a tensor is quasi-positive if it is nonngeative everywhere and positive definite at some point. Actually Theorem 5.8 follows from Theorem 1.1 and the following result.
Proposition 5.9.
Let $E$ be a holomorphic vector bundle over a compact Kähler manifold $(M,\omega_{g})$. Suppose that there exists a smooth Hermitian metric $h_{0}$ on $E$ such that the Hermitian-Yang-Mills tensor $\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}$ is quasi-positive definite. Then there exists a smooth Hermitian metric $h_{1}$ such that $\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{1}}$ is positive definite.
Proof.
We employ the strategy utilized in the proof of Theorem 5.6. We fix the center point $h_{0}\in\mathrm{Herm}^{+}(E)$ and so $F(h_{0})=\mathrm{Id}_{E}$. We show that $G$ is invertible around $h_{0}$. In particular, when $\varepsilon>0$ is small, there exists some $h_{1}$ on $E$ such that
(5.84) $\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{1}}=\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}+\varepsilon h_{0}>0.$
Indeed, for any $\Psi\in\mathrm{Herm}(E,h_{0})$,
(5.85) $\mathscr{L}_{\mathrm{Id}_{E}}(\Psi)=\Delta_{{\partial}^{h_{0}}}\Psi+\Omega_{0}\cdot\Psi.$
where $\Omega_{0}=S^{h_{0}}\in\Gamma(M,E^{*}\otimes E)$. To apply the framework established in Theorem 5.6, it suffices to demonstrate the existence of a constant $C=C(M,\omega_{g},h_{0})$ such that for each $\Psi\in W^{1,2}(M,E^{*}\otimes E)$,
(5.86) $(\mathscr{L}_{\mathrm{Id}_{E}}\Psi,\Psi)\geq C(\Psi,\Psi),$
and this implies
(5.87) $(\mathscr{L}_{\mathrm{Id}_{E}}\Psi,\Psi)\geq\min\left\{\frac{C}{2},\frac{1}{2}\right\}\|\Psi\|_{W^{1,2}(M,E^{*}\otimes E)}.$
If it is not valid, for any $m\in\mathbb{N}$, there exists nonzero $\Psi_{m}\in W^{1,2}(M,E^{*}\otimes E)$ such that
(5.88) $({\partial}^{h_{0}}\Psi_{m},{\partial}^{h_{0}}\Psi_{m})+(\Omega\cdot\Psi_{m},\Psi_{m})<\frac{1}{m}(\Psi_{m},\Psi_{m}).$
We use the normalization trick and define
(5.89) $\Phi_{m}:=\frac{\Psi_{m}}{\left\|\Psi_{m}\right\|_{L^{2}(M,E^{*}\otimes E)}}.$
One has $\left\|\Phi_{m}\right\|_{L^{2}(M,E^{*}\otimes E)}=1$ and
(5.90) $({\partial}^{h_{0}}\Phi_{m},{\partial}^{h_{0}}\Phi_{m})+(\Omega\cdot\Phi_{m},\Phi_{m})<\frac{1}{m}(\Phi_{m},\Phi_{m})=\frac{1}{m}.$
By (5.90) and the normalization condition, there exists $C=C(M,\omega_{g})$ such that,
(5.91) $\left\|\Phi_{m}\right\|_{W^{1,2}(M,E^{*}\otimes E)}\leq C.$
By the weak compactness theorem and the compact embedding of $W^{1,2}(M,E^{*}\otimes E)\subset L^{2}(M,E^{*}\otimes E)$, there exists a subsequence $\{\Phi_{m_{k}}\}$ and $\Phi\in W^{1,2}(M,E^{*}\otimes E)$ such that $\Phi_{m_{k}}\rightarrow\Phi$ in the weak $W^{1,2}(M,E^{*}\otimes E)$ sense and $\Phi_{m_{k}}\rightarrow\Phi$ in $L^{2}(M,E^{*}\otimes E)$ sense. It is clear that $\left\|\Phi\right\|_{L^{2}(M,E^{*}\otimes E)}=1$. By the weak convergence theorem,
[…]
Therefore, we obtain that ${\partial}^{h_{0}}\Phi=0$. On the other hand, by the $L^{2}$-convergence,
(5.93) $(\Omega\cdot\Phi,\Phi)=\lim_{m_{k}\rightarrow\infty}(\Omega\cdot\Phi_{m_{k}},\Phi_{m_{k}})\leq\lim_{m_{k}\rightarrow\infty}\frac{1}{m_{k}}=0.$
Since $\Omega$ is quasi-positive, there exists an open set $U\subset M$ such that $\Phi\equiv 0$ on $U$. Since ${\partial}^{h_{0}}\Phi=0$, $\Phi\equiv 0$ on $M$ which contradicts to the fact $\left\|\Phi\right\|_{L^{2}(M,E^{*}\otimes E)}=1$. ∎
Remark 5.10.
Proposition 5.9 provides an alternative proof of the main results in [Yang24] and [LL22].
6. Chern number inequalities
In this section, we prove Theorem 1.6 in a slightly general setting:
Theorem 6.1.
Let $(M,\omega_{g})$ be a compact Kähler manifold and $E$ be a holomorphic vector bundle of rank $r$ over $M$. Suppose that there exists a Hermitian metric $h_{0}$ on $E$ satisfying
(6.1) $a\cdot h_{0}\leq\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}\leq b\cdot h_{0},$
for some constants $a,b\in{\mathbb{R}}$. Then the following Chern number inequality holds
(6.2) $\int_{M}\left((r-1)c^{2}_{1}(E)-2rc_{2}(E)\right)\wedge\omega_{g}^{n-2}\leq\frac{r(r-1)\left(b-a\right)^{2}}{8\pi^{2}n^{2}}\int_{M}\omega_{g}^{n},$
if $ab>0$ or there exists a holomorphic line bundle $L$ on $M$ satisfying $\int_{M}c_{1}(L)\wedge\omega_{g}^{n}\neq 0$.
Proof.
We assume that $b\geq a>0$. By Theorem 1.1, there exists a smooth Hermitian metric $h$ on $E$ such that
(6.3) $\Lambda_{\omega_{g}}\sqrt{-1}R^{h}=h_{0}.$
In the following, we work on the abstract Hermitian vector bundle $(E,h)$. Let $\{z^{i}\}$ be local holomorphic coordinates on $M$ and $\{e_{\alpha}\}$ be holomorphic frames of $E$. The Chern curvature tensor $R^{h}\in\Gamma(M,\Lambda^{1,1}T^{*}M\otimes E^{*}\otimes\overline{E}^{*})$ is given by
$R^{h}=R^{h}_{i\overline{j}\alpha\overline{\beta}}dz^{i}\wedge d\overline{z}^{j}\otimes e^{\alpha}\otimes\overline{e}^{\beta}.$
We set
$R^{(1)}_{i\overline{j}}=h^{\alpha\overline{\beta}}R^{h}_{i\overline{j}\alpha\overline{\beta}},\quad R^{(2)}_{\alpha\overline{\beta}}=g^{i\overline{j}}R^{h}_{i\overline{j}\alpha\overline{\beta}}\quad\mbox{and}\quad s_{h}=g^{i\overline{j}}R^{(1)}_{i\overline{j}}$.
The following expressions are well-known:
(6.4) $\int_{M}c^{2}_{1}(E)\wedge\omega_{g}^{n-2}=\frac{1}{4\pi^{2}n(n-1)}\int_{M}\left(s_{h}^{2}-\left|\mathrm{Ric}^{(1)}\right|_{g}^{2}\right)\omega_{g}^{n},$
and
(6.5) $\int_{M}c_{2}(E)\wedge\omega_{g}^{n-2}=\frac{1}{8\pi^{2}n(n-1)}\int_{M}\left(s_{h}^{2}-\left|\mathrm{Ric}^{(1)}\right|_{g}^{2}-\left|\mathrm{Ric}^{(2)}\right|^{2}_{h}+|R^{h}|_{g,h}^{2}\right)\omega_{g}^{n}.$
We define $T\in\Gamma(M,\Lambda^{1,1}T^{*}M\otimes E^{*}\otimes\overline{E}^{*})$:
(6.6) $T_{i\overline{j}\alpha\overline{\beta}}=R^{h}_{i\overline{j}\alpha\overline{\beta}}-\frac{1}{n}g_{i\overline{j}}R^{(2)}_{\alpha\overline{\beta}}-\frac{1}{r}R^{(1)}_{i\overline{j}}h_{\alpha\overline{\beta}}+\frac{1}{nr}g_{i\overline{j}}h_{\alpha\overline{\beta}}s_{h}.$
A straightforward computation shows that
(6.7) $|T|^{2}=|R^{h}|^{2}-\frac{1}{n}\left|\mathrm{Ric}^{(2)}\right|^{2}-\frac{1}{r}\left|\mathrm{Ric}^{(1)}\right|^{2}+\frac{1}{nr}s_{h}^{2}.$
By using $|T|^{2}$, one has
$\int_{M}\left(2rc_{2}(E)-(r-1)c^{2}_{1}(E)\right)\wedge\omega_{g}^{n-2}$
$=$ $\frac{1}{4\pi^{2}n(n-1)}\int_{M}\left(s_{h}^{2}-\left|\mathrm{Ric}^{(1)}\right|^{2}-r\left|\mathrm{Ric}^{(2)}\right|^{2}+r|R^{h}|^{2}\right)\omega_{g}^{n}$
$=$ $\frac{1}{4\pi^{2}n(n-1)}\int_{M}\left(r|T|^{2}+\frac{n-1}{n}\left(s_{h}^{2}-r\left|\mathrm{Ric}^{(2)}\right|^{2}\right)\right)\omega_{g}^{n}.$
Since $\Lambda_{\omega_{g}}\sqrt{-1}R^{h}=h_{0}$, we get
(6.8) $s_{h}=\mathrm{tr}_{h}h_{0},\quad\mbox{ and}\quad\left|\mathrm{Ric}^{(2)}\right|^{2}=|h_{0}|_{h}^{2}.$
Therefore, we obtain
$\int_{M}\left(2rc_{2}(E)-(r-1)c^{2}_{1}(E)\right)\wedge\omega_{g}^{n-2}$
$=$ $\frac{1}{4\pi^{2}n(n-1)}\int_{M}\left(r|T|^{2}+\frac{n-1}{n}\left((\mathrm{tr}_{h}h_{0})^{2}-r|h_{0}|^{2}_{h}\right)\right)\omega_{g}^{n}.$
Since $0<a\cdot h_{0}\leq\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}\leq b\cdot h_{0}$ and $\Lambda_{\omega_{g}}\sqrt{-1}R^{h}=h_{0}$, we deduce that
(6.9) $0<a\cdot\sqrt{-1}\Lambda_{\omega_{g}}R^{h}\leq\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}\leq b\cdot\sqrt{-1}\Lambda_{\omega_{g}}R^{h}.$
By Theorem 1.2, we obtain
(6.10) $a\cdot h\leq h_{0}\leq b\cdot h.$
In particular, the following estimates hold
(6.11) $(\mathrm{tr}_{h}h_{0})^{2}-r|h_{0}|^{2}_{h}\geq-\frac{r(r-1)}{2}(b-a)^{2}\geq-r^{2}(b^{2}-a^{2}).$
Finally, we establish the Chern number inequality
$\int_{M}\left(2rc_{2}(E)-(r-1)c^{2}_{1}(E)\right)\wedge\omega_{g}^{n-2}\geq-\frac{r(r-1)\left(b-a\right)^{2}}{8\pi^{2}n^{2}}\int_{M}\omega_{g}^{n}+\frac{r}{4\pi^{2}n(n-1)}\int_{M}|T|^{2}\omega_{g}^{n}.$
For the case $a\leq b<0$, we know that
(6.12) $0<-b\cdot h_{0}^{*}\leq\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}^{*}}\leq-a\cdot h_{0}^{*}$
as Hermitian tensors in $\Gamma(M,E\otimes\overline{E})$ under the isomorphism $E\cong(E^{*})^{*}$. Hence, there exists a Hermitian metric $h^{*}$ on $E^{*}$ such that
(6.13) $\Lambda_{\omega_{g}}\sqrt{-1}R^{h^{*}}=h^{*}_{0}.$
Moreover, one has
(6.14) $0<-b\cdot\sqrt{-1}\Lambda_{\omega_{g}}R^{h^{*}}\leq-\Lambda_{\omega_{g}}\sqrt{-1}R^{h^{*}_{0}}\leq-a\cdot\sqrt{-1}\Lambda_{\omega_{g}}R^{h^{*}},$
and so $bh^{*}\leq h_{0}^{*}\leq-ah^{*}$. This implies
(6.15) $\int_{M}\left(2rc_{2}(E^{*})-(r-1)c^{2}_{1}(E^{*})\right)\wedge\omega_{g}^{n-2}\geq-\frac{r(r-1)\left(b-a\right)^{2}}{8\pi^{2}n^{2}}\int_{M}\omega_{g}^{n}.$
Since $c_{1}^{2}(E^{*})=c_{1}^{2}(E)$ and $c_{2}(E^{*})=c_{2}(E)$, we get the result.
Suppose that there exists a holomorphic line bundle $L$ over $M$ with $\int_{M}c_{1}(L)\wedge\omega_{g}^{n-1}>0$. It is easy to show that there exists a smooth Hermitian metric $h^{L}$ on $L$ such that
(6.16) $\Lambda_{\omega_{g}}\sqrt{-1}R^{h^{L}}=c_{0}=\frac{n\int_{M}c_{1}(L)\wedge\omega_{g}^{n-1}}{\int_{M}\omega_{g}^{n}}>0.$
We consider the twisted Hermitian vector bundle $E_{k}:=E\otimes L^{k}$. The Hermitian-Yang-Mills tensor of $(E_{k},h_{0}\otimes h^{kL})$ is given by
(6.17) $\Lambda_{\omega_{g}}\left(\sqrt{-1}R^{h_{0}\otimes h^{kL}}\right)=\left(\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}\right)\otimes h^{kL}+h_{0}\otimes\Lambda_{\omega_{g}}\left(\sqrt{-1}R^{h^{kL}}\right).$
In particular,
(6.18) $\Lambda_{\omega_{g}}\left(\sqrt{-1}R^{h_{0}\otimes h^{kL}}\right)=\left(\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}+kc_{0}h_{0}\right)\otimes h^{kL}.$
We fix a positive integer $k$ such that $a+kc_{0}>0$. Then the Hermitian-Yang-Mills tensor of $(E_{k},h_{0}\otimes h^{kL})$ satisfies
(6.19) $0<(a+kc_{0})(h_{0}\otimes h^{kL})\leq\Lambda_{\omega_{g}}\left(\sqrt{-1}R^{h_{0}\otimes h^{kL}}\right)\leq(b+kc_{0})(h_{0}\otimes h^{kL}).$
In particular, by the first case, we deduce that
(6.20) $\int_{M}\left(2rc_{2}(E\otimes L^{k})-(r-1)c^{2}_{1}(E\otimes L^{k})\right)\wedge\omega_{g}^{n-2}\geq-\frac{r(r-1)\left(b-a\right)^{2}}{8\pi^{2}n^{2}}\int_{M}\omega_{g}^{n}.$
On the other hand, it is well-known that
$\int_{M}\left(2rc_{2}(E\otimes L^{k})-(r-1)c^{2}_{1}(E\otimes L^{k})\right)\wedge\omega_{g}^{n-2}=\int_{M}\left(2rc_{2}(E)-(r-1)c^{2}_{1}(E)\right)\wedge\omega_{g}^{n-2}.$
Hence, we obtain the result. ∎
Remark 6.2.
It is expected that the projectivity condition in Theorem 1.6 is redundant, i.e., the Chern number inequality (1.12) holds for any $a,b\in{\mathbb{R}}$ satisfying (1.11).
Corollary 6.3.
Let $M$ be a Fano manifold. If there exists a Kähler metric $\omega_{g}$ such that
(6.21) $a\omega_{g}\leq\mathrm{Ric}(\omega_{g})\leq b\omega_{g}$
for some constants $a>0,b>0$, then the following Chern number inequality holds
(6.22) $\int_{M}\left((n-1)c^{2}_{1}(M)-2nc_{2}(M)\right)\wedge\omega_{g}^{n-2}\leq\frac{(n-1)\left(b-a\right)^{2}}{8\pi^{2}n}\int_{M}\omega_{g}^{n}.$
Remark 6.4.
In the proof of Corollary 6.3, we use the existence of a Hermitian metric $h$ on $M$ satisfying
(6.23) $\Lambda_{\omega_{g}}\sqrt{-1}R^{h}=g,$
where $g$ is a Kähler metric. This condition enriches the geometric interplay between $h$ and $g$, allowing for sharper control in the curvature estimates. In particular, when $h$ is a Kähler or balanced metric– i.e., $d\omega^{n-1}_{h}=0$–the associated Chern number inequality can be strengthened to
(6.24) $\int_{M}\left(nc^{2}_{1}(M)-2(n+1)c_{2}(M)\right)\wedge\omega_{g}^{n-2}\leq\frac{(n^{2}-1)\left(b-a\right)^{2}}{8\pi^{2}n^{2}}\int_{M}\omega_{g}^{n}.$
This refines and extends Yau’s classical Chern number inequalities [Yau77] for Fano Kähler-Einstein manifolds.
7. The prescribed Hermitian-Yang-Mills tensor equation on line bundles
In this section, we investigate the prescribed Hermitian-Yang-Mills equation on line bundles and clarify its connection with RC-positivity framework. Moreover, we show that the condition $P>0$ in Theorem 1.1 is necessary:
Proposition 7.1.
Let $L$ be a holomorphic line bundle over a compact Kähler manifold $(M,\omega_{g})$. Suppose that there exists a Hermitian metric $h_{0}$ on $L$ such that $\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}>0$. If $P\in\Gamma(M,L^{*}\otimes\overline{L}^{*})$ fails to be Hermitian positive definite, then the equation
(7.1) $\Lambda_{\omega_{g}}\sqrt{-1}R^{h}=P$
may admit two distinct smooth solutions.
Lemma 7.2.
Let $L$ be a holomorphic line bundle over a compact Kähler manifold $(M,\omega_{g})$. Then for any Hermitian metric $h_{0}$ on $L$,
(7.2) $S^{h_{0}}=\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}=s_{h_{0}}\cdot h_{0}\in\Gamma(M,L^{*}\otimes\overline{L}^{*})$
where $s_{h_{0}}$ is the Chern scalar curvature of $(L,h_{0})$.
Proof.
It is well-known for any Hermitian metrics $h_{1},h_{2}\in\Gamma(M,L^{*}\otimes\overline{L}^{*})$, there exists some smooth function $\varphi\in C^{\infty}(M,{\mathbb{R}})$ such that
(7.3) $h_{2}=e^{-\varphi}h_{1}.$
Moreover, if $P\in\Gamma(M,L^{*}\otimes\overline{L}^{*})$ is a Hermitian tensor, then $P=fh_{1}$ some $f\in C^{\infty}(M,{\mathbb{R}})$. Let $e$ be a local holomorphic base of $L$ and $h_{0}=h_{0}(e,e)$. By using the curvature formula, one has
(7.4) $S^{h_{0}}=\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}=-g^{i\overline{j}}\frac{{\partial}^{2}\log h_{0}}{{\partial}z^{i}{\partial}\overline{z}^{j}}\cdot h_{0}e^{*}\otimes\overline{e}^{*}=-\left(\Delta_{g}\log h_{0}\right)h_{0}e^{*}\otimes\overline{e}^{*}.$
The Chern scalar curvature function is
(7.5) $s_{h_{0}}=\mathrm{tr}_{h_{0}}S^{h_{0}}=-\Delta_{g}\log h_{0}.$
Hence, $S^{h_{0}}=\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}=s_{h_{0}}\cdot h_{0}$
Proposition 7.3.
Let $(L,h_{0})$ be a Hermitian holomorphic line bundle over a compact Kähler manifold $(M,\omega_{g})$ and $S^{h_{0}}=Fh_{0}$ for some $F\in C^{\infty}(M,{\mathbb{R}})$. Suppose that $P=Gh_{0}\in\Gamma(M,L^{*}\otimes\overline{L}^{*})$ for some $G\in C^{\infty}(M,{\mathbb{R}})$. Then there exists some Hermitian metric $h$ on $L$ satisfying
(7.6) $\Lambda_{\omega_{g}}\sqrt{-1}R^{h}=P,$
if and only if there exists some $\varphi\in C^{\infty}(M,{\mathbb{R}})$ such that
(7.7) $\left(\Delta_{g}\varphi+F\right)e^{-\varphi}=G.$
In this correspondence, one has $h=e^{-\varphi}h_{0}$.
Proof.
It is known that $S^{h_{0}}=Fh_{0}$ where
(7.8) $F=-\Delta_{g}\log h_{0}.$
Hence, if $S^{h}=P=Gh_{0}$ for some Hermitian metric $h$, then
(7.9) $G=-\Delta_{g}\log h\cdot\left(h\cdot h_{0}^{-1}\right).$
On the other hand, $h=e^{-\varphi}h_{0}$ for some $\varphi\in C^{\infty}(M,{\mathbb{R}})$. Therefore,
$\Delta_{g}\log h=-\Delta_{g}\log h_{0}+\Delta_{g}\varphi$
and so $G=\left(\Delta_{g}\varphi+F\right)e^{-\varphi}$. ∎
The following result is a consequence of Theorem 1.1 and Proposition 7.3, which is a line bundle version of the Kazdan-Warner theorem ([KW74]).
Corollary 7.4.
Let $L$ be a holomorphic line bundle over a compact Kähler manifold $(M,\omega_{g})$. Suppose that there exists a Hermitian metric $h_{0}$ on $L$ such that its Chern scalar curvature $F=s_{h_{0}}>0$, then for any positive smooth function $G$, the following equation has the unique smooth solution $\varphi$:
(7.10) $e^{-\varphi}(F+\Delta_{g}\varphi)=G.$
For a detailed discussion on the existence of Hermitian metrics with positive Chern scalar curvature, positive total Chern scalar curvature and RC-positivity for line bundles, we refer to [Yang18], [Yang19a] and [Yang19b]. This is also a motivation of this project for vector bundles. In this following, we show that the equation (7.10) may admit multiple solutions when $G$ is not positive.
Proof of Proposition 7.1.
Let $L$ be an ample line bundle or a line bundle with
$\int_{M}c_{1}(L)\wedge\omega_{g}^{n-1}>0$
over a projective Kähler manifold $(M,\omega_{g})$. There exists a Hermitian metric $h_{0}$ on $L$ such that $S^{h_{0}}=\Lambda_{\omega_{g}}\sqrt{-1}R^{h}=Fh_{0}>0$ ([Yang19a]). Recall that
(7.11) $F=-\Delta_{g}\log h_{0}=s_{h_{0}}.$
Let $f\in C^{\infty}(M,\mathbb{R})$ be a smooth function satisfying
  1. (1)
    $f\geq 0$ and there exists $p\in M$ such that $f(p)=0$ and $f(x)>0$ if $x\neq p$.
  2. (2)
    There exists some small $\delta>0$ and $q\in B(p,\delta)$ such that
    (7.12) $f(q)=\begin{cases}e^{-1/r^{n}}&\text{if }q\neq p\\ 0&\text{if }q=p,\end{cases}$
    where $r=d_{g}(p,q)$ is the distance function.
We consider the map $Q:(0,+\infty)\rightarrow\mathbb{R}$ defined by
(7.13) $Q(t)=\int_{M}\left(\Delta_{g}f\right)/(e^{f+t}-1)\cdot\omega_{g}^{n}.$
We claim that
(7.14) $\lim_{t\rightarrow 0^{+}}Q(t)=+\infty,\quad\lim_{t\rightarrow+\infty}Q(t)=0.$
Indeed, by Stokes’ theorem, one has
$2Q(t)$ $=$ $2\int_{M}\left(\Delta_{g}f\right)/(e^{f+t}-1)\cdot\omega_{g}^{n}=-\int_{M}g\left(\nabla f,\nabla(e^{f+t}-1)^{-1}\right)\cdot\omega_{g}^{n}$
$=$ $\int_{M}|\nabla f|^{2}\cdot(e^{f+t}-1)^{-2}\cdot e^{f+t}\cdot\omega_{g}^{n}=\int_{M}|\nabla f|^{2}\cdot\left(e^{(f+t)}+e^{-(f+t)}-2\right)^{-1}\cdot\omega_{g}^{n}.$
Therefore, $Q$ is a decreasing in $t$ and by monotone convergence theorem, one has
(7.15) $\quad\lim_{t\rightarrow+\infty}Q(t)=0.$
On the other hand, by monotone convergence theorem
(7.16) $\lim_{t\rightarrow 0^{+}}Q(t)=\int_{M\setminus\{p\}}|\nabla f|^{2}\cdot(e^{f}+e^{-f}-2)^{-1}\cdot\omega_{g}^{n}.$
Moreover, there is a constant $A=A(M,\delta)>0$ such that for any $q\in B(p,\delta)$, one has
(7.17) $e^{f}+e^{-f}-2\leq Af^{2},$
and so
$\lim_{t\rightarrow 0^{+}}Q(t)$ $\geq$ $A^{-1}\int_{M\setminus\{p\}}|\nabla f|^{2}/f^{2}\cdot\omega_{g}^{n}=A^{-1}\int_{B(p,\delta)\setminus\{p\}}|\nabla\mathrm{log}(f)|^{2}\cdot\omega_{g}^{n}$
$=$ $n^{2}A^{-1}\int_{B(p,\delta)\setminus\{p\}}r^{-(2n+2)}\cdot\omega_{g}^{n}=+\infty.$
By continuity of $Q(t)$, there exists some $t_{0}\in(0,+\infty)$ such that
(7.18) $\int_{M}\left(\Delta_{\omega_{g}}f\right)/\left(e^{f+t_{0}}-1\right)\cdot\omega_{g}^{n}=n\int_{M}c_{1}(L)\wedge\omega_{g}^{n-1}>0.$
Let $\Phi=f+t_{0}$. Then one has
(7.19) $\int_{M}\left(\Delta_{g}\Phi\right)/\left(e^{\Phi}-1\right)\cdot\omega_{g}^{n}=n\int_{M}c_{1}(L)\wedge\omega_{g}^{n-1}=\int_{M}s_{h_{0}}\cdot\omega_{g}^{n}=\int_{M}F\cdot\omega_{g}^{n}.$
By Hodge theory, there exists a smooth function $\psi\in C^{\infty}(M,\mathbb{R})$ such that
(7.20) $\frac{\Delta_{g}\Phi}{e^{\Phi}-1}=F-\Delta_{g}\psi.$
If we define
(7.21) $G:=\frac{\Delta_{g}\Phi}{e^{\Phi}-1}\cdot e^{\psi}$
then it is obvious that the equation
(7.22) $e^{-\varphi}(F+\Delta_{g}\varphi)=G$
has two different solutions $\varphi=-\psi$ and $\varphi=\Phi-\psi$. In particular, the equation
(7.23) $\Lambda_{\omega_{g}}\sqrt{-1}R^{h}=Gh_{0}$
has two different solutions $h_{1}=e^{\psi}\cdot h_{0}$ and $h_{2}=e^{\psi-\Phi}h_{0}$. ∎
Example 7.5.
Let $(L,h_{0})$ be a Hermitian holomorphic trivial line bundle over a compact Kähler manifold $(M,\omega_{g})$. Suppose $P=Gh_{0}\in\Gamma(M,L^{*}\otimes\overline{L}^{*})$ for some $G\in C^{\infty}(M,{\mathbb{R}})$ satisfying
$\int_{M}G\omega_{g}^{n}<0$
then the equation $\Lambda_{\omega_{g}}\sqrt{-1}R^{h}=P$ has no solution. Indeed, let $\varphi\in C^{\infty}(M,{\mathbb{R}})$ and $h=e^{-\varphi}h_{0}$. One has
(7.24) $S^{h}=\Lambda_{\omega_{g}}\sqrt{-1}R^{h}=\left(e^{-\varphi}\Delta_{\omega_{g}}\varphi\right)\cdot h_{0}.$
Moreover,
(7.25) $\int_{M}\mathrm{tr}_{h_{0}}S^{h}\cdot\omega_{g}^{n}=\int_{M}\left(e^{-\varphi}\Delta_{\omega_{g}}\varphi\right)\cdot\omega_{g}^{n}=\int_{M}e^{-\varphi}|\nabla\varphi|^{2}\omega_{g}^{n}\geq 0.$
In particular, if there exists some Hermitian metric $h$ on $L$ such that $\Lambda_{\omega_{g}}\sqrt{-1}R^{h}=P$, then $\int_{M}G\omega_{g}^{n}\geq 0$. This is a contradiction.